### v1

# Refined canonical stable Grothendieck polynomials and their duals, Part 2

Abstract.
This paper is the sequel of the paper under the same title with part 1, where we introduced refined canonical stable Grothendieck polynomials and their duals with two families of infinite parameters. In this paper we give combinatorial interpretations for these polynomials using generalizations of set-valued tableaux and reverse plane partitions, respectively. Our results extend to their flagged and skew versions.
1. Introduction
This paper is a continuation of the study of refined canonical stable Grothendieck polynomials $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ and their duals $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ introduced in the previous paper [10]. The main goal of this paper is to give combinatorial models for $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ using generalizations of set-valued tableaux and reverse plane partitions, respectively. In doing so, we also generalize $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ to their flagged and skew versions.
We briefly review known results on Grothendieck polynomials relevant to the current paper. For a more detailed history of Grothendieck polynomials, see the introduction of the previous paper [10].
In 1982, Lascoux and Schützenberger [14] introduced Grothendieck polynomials, which are $K$-theoretic analogues of Schubert polynomials. Fomin and Kirillov [5] added a parameter $\beta$ to Grothendieck polynomials and studied stable limits of $\beta$-Grothendieck polynomials. The stable Grothendieck polynomials $G_{\lambda}(\bm{x})$ of Grassmannian type are indexed by partitions and form a basis of (a completion of) the symmetric function space.
It is a classical result that the Schur function $s_{\lambda}(\bm{x})$ has a combinatorial model using semistandard Young tableaux. In 2002, Buch [2] found a similar combinatorial model for $G_{\lambda/\mu}(\bm{x})$ by introducing set-valued tableaux. The dual Grothendieck polynomials $g_{\lambda}(\bm{x})$ are dual to $G_{\lambda}(\bm{x})$ with respect to the Hall inner product. In 2007, Lam–Pylyavskyy [13] found a combinatorial model for $g_{\lambda/\mu}(\bm{x})$ using reverse plane partitions. Since then, Grothendieck polynomials $G_{\lambda}(\bm{x})$ and their duals $g_{\lambda}(\bm{x})$ have been shown to have many interesting combinatorial properties and generalizations.
Yeliussizov [20] introduced canonical stable Grothendieck polynomials $G^{(\alpha,\beta)}_{\lambda}(\bm{x})$ and their duals $g^{(\alpha,\beta)}_{\lambda}(\bm{x})$ with two parameters $\alpha$ and $\beta$, which behave nicely under the $\omega$ involution on the symmetric function space. Motivated by their study of Brill–Noether varieties in [4], Chan–Pflueger [3] introduced Grothendieck polynomials $RG_{\lambda}(\bm{x};\bm{\beta})$ with a family $\bm{\beta}=(\beta_{1},\beta_{2},\dots)$ of infinite parameters. In a completely different context, Galashin–Grinberg–Liu [6] introduced dual Grothendieck polynomials $\tilde{g}_{\lambda}(\bm{x};\bm{\beta})$ with infinite parameters $\bm{\beta}$, which turn out to be dual to $RG_{\lambda}(\bm{x};\bm{\beta})$ with respect to the Hall inner product.
In the previous paper [10] we introduced refined canonical stable Grothendieck polynomials $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ and their duals $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ with two families $\bm{\alpha}$ and $\bm{\beta}$ of infinite parameters, which generalize all variations of Grothendieck polynomials mentioned so far. In this paper we find combinatorial models for $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ using generalizations of set-valued tableaux and reverse plane partitions, respectively. Our results, in fact, apply to further generalizations of $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ described below.
There is another way to generalize Grothendieck polynomials by imposing “flag conditions” on their tableau models, where a flag condition means that the entries in a row or column are bounded by specific numbers. Gessel and Viennot [7] found a Jacobi–Trudi-like formula for flagged Schur functions, which are generating functions for flagged semistandard Young tableaux. Wachs [19] gave another proof of this result using natural recursions coming from flagged semistandard Young tableaux. Knutson, Miller, and Yong [12] showed that Grothendieck polynomials associated with vexillary permutations are generating functions for flagged set-valued tableaux. Matsumura [17] introduced flagged Grothendieck polynomials $G_{\lambda/\mu,f/g}(\bm{x})$ as generating functions for flagged set-valued tableaux and proved their Jacobi–Trudi-like formula. For the dual case, Grinberg [8] first considered flagged versions $\tilde{g}^{\operatorname{row}(\bm{r},\bm{s})}_{\lambda/\mu}(\bm{x};\bm{\beta})$ of $\tilde{g}_{\lambda/\mu}(\bm{x};\bm{\beta})$, which were further studied by Kim [11].
[…]
Variations of Grothendieck polynomials introduced in how to specialize
$G_{\nu/\lambda}(\bm{x})$ Buch [2] $G_{\nu/\lambda}(\bm{x};\bm{0},\bm{1})$
$g_{\lambda/\mu}(\bm{x})$ Lam–Pylyavskyy [13] $g_{\lambda/\mu}(\bm{x};\bm{0},\bm{1})$
$G^{(\alpha,\beta)}_{\lambda}(\bm{x})$ Yeliussizov [20] $G_{\lambda}(\bm{x};\bm{\alpha}_{0},-\bm{\beta}_{0})$
$g^{(\alpha,\beta)}_{\lambda}(\bm{x})$ Yeliussizov [20] $g_{\lambda}(\bm{x};-\bm{\alpha}_{0},\bm{\beta}_{0})$
$RG_{\lambda}(\bm{x};\bm{\beta})$ Chan–Pflueger [3] $G_{\lambda}(\bm{x};\bm{0},\bm{\beta})$
$\tilde{g}_{\lambda/\mu}(\bm{x};\bm{\beta})$ Galashin–Grinberg–Liu [6] $g_{\lambda/\mu}(\bm{x};\bm{0},\bm{\beta})$
$G_{\lambda/\mu,f/g}(\bm{x})$ Matsumura [17] $G^{\operatorname{row}(g,f)}_{\lambda/\mu}(\bm{x};\bm{0},-\bm{\beta}_{0})$
$\tilde{g}^{\operatorname{row}(\bm{r},\bm{s})}_{\lambda/\mu}(\bm{x};\bm{\beta})$ Grinberg [8], Kim [11] $g^{\operatorname{row}(\bm{r},\bm{s})}_{\lambda/\mu}(\bm{x};\bm{0},\bm{\beta})$
In what follows we illustrate the main results in this paper. To do this we recall the definitions of $G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ in our previous paper [10] and their Jacobi–Trudi-like formulas; see Section 2 for the definitions of undefined terms.
Definition 1.1.
Let $\lambda$ be a partition $\lambda$ with at most $n$ parts. The refined canonical stable Grothendieck polynomial $G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ and the refined dual canonical stable Grothendieck polynomial $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ are defined by
(1.1) $G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ $=\frac{\det\left(x_{j}^{\lambda_{i}+n-i}\dfrac{(1-\beta_{1}x_{j})\cdots(1-\beta_{i-1}x_{j})}{(1-\alpha_{1}x_{j})\cdots(1-\alpha_{\lambda_{i}}x_{j})}\right)_{1\leq i,j\leq n}}{\prod_{1\leq i<j\leq n}(x_{i}-x_{j})},$
(1.2) $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ $=\frac{\det(h_{\lambda_{i}+n-i}[x_{j}-A_{\lambda_{i}-1}+B_{i-1}])_{1\leq i,j\leq n}}{\prod_{1\leq i<j\leq n}(x_{i}-x_{j})}.$
In the first part [10] we proved that $G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ have Jacobi-Trudi-like formulas.
Theorem 1.2 ([10]).
For a partition $\lambda$ with at most $n$ parts, we have
(1.3) $G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ $=C_{n}\det\left(h_{\lambda_{i}-i+j}[X_{n}\ominus(A_{\lambda_{i}}-B_{i-1}+B_{j})]\right)_{i,j=1}^{n},$
(1.4) $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ $=\det\left(h_{\lambda_{i}-i+j}[X_{n}-A_{\lambda_{i}-1}+B_{i-1}-B_{j-1}]\right)_{i,j=1}^{n},$
where $C_{n}=\prod_{i,j=1}^{n}(1-\beta_{i}x_{j})$ and $h_{n}[Y\ominus Z]=\sum_{a-b=n}h_{a}[Y]h_{b}[Z]$.
The main results in this paper are the following combinatorial interpretations for $G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ and $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ using “marked multiset-valued tableaux” and “marked reverse plane partitions”, which are generalizations of set-valued tableaux and reverse plane partitions, respectively; see Sections 3 and 4 for their precise definitions.
Theorem 1.3.
For a partition $\lambda$, we have
(1.5) $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\sum_{T\in\operatorname{MMSVT}(\lambda)}\operatorname{wt}(T),$
(1.6) $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\sum_{T\in{\operatorname{MRPP}}(\lambda)}\operatorname{wt}(T),$
where $\operatorname{MMSVT}(\lambda)$ and ${\operatorname{MRPP}}(\lambda)$ are the sets of marked multiset-valued tableaux and marked reverse plane partitions of shape $\lambda$, respectively.
Our strategy to prove Theorem 1.3 is as follows. Instead of proving (1.5) and (1.6) directly, we consider their flagged versions $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ for which we can use inductive arguments in the same vein as in Wachs [19]. Using computer experiments, we were able to find suitable Jacobi–Trudi-like formulas for $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ that generalize (1.3) and (1.4), respectively; see Theorem 1.4 below. Thanks to the flag conditions, $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ have natural recurrences. We then prove the Jacobi–Trudi-like formulas by showing that the determinants also have the same recurrences. Although our proofs require many technical lemmas, they are elementary and we need nothing more complex than basic techniques in linear algebra such as row operations.
Our methods also work for column-flagged versions. In summary we prove the following Jacobi–Trudi-like formulas. Note that (1.7) and (1.9) generalize (1.3) and (1.4), hence imply (1.5) and (1.6), respectively. We also note that (1.8) is proved under a weaker condition; see Theorem 3.8.
Theorem 1.4.
Let $\lambda$ and $\mu$ be partitions with at most $n$ parts and $\mu\subseteq\lambda$. Let $\bm{r}=(r_{1},\dots,r_{n})$ and $\bm{s}=(s_{1},\dots,s_{n})$ be positive integer sequences of length $n$. Suppose that $r_{i}\leq r_{i+1}$ and $s_{i}\leq s_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$. Then
(1.7) $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=C\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j})]\right)_{i,j=1}^{n},$
(1.8) $G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=D\det\left(e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{i-1}-A_{j}-B_{\lambda_{i}}+B_{\mu_{j}})]\right)_{i,j=1}^{n},$
(1.9) $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{\lambda_{i}-1}+A_{\mu_{j}}+B_{i-1}-B_{j-1}]\right)_{i,j=1}^{n},$
(1.10) $g_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\det\left(e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1}-B_{\mu_{j}}]\right)_{i,j=1}^{n},$
where
$C=\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l}),\qquad D=\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\alpha_{i}x_{l})^{-1}.$
The rest of this paper is organized as follows. In Section 2 we recall necessary definitions and lemmas. In Section 3 we prove (1.7) and (1.8). We also show that our result implies the Jacobi–Trudi-like formula for $G_{\lambda/\mu,f/g}(\bm{x})$ due to Matsumura [17]. In Section 4 we prove (1.9) and (1.10).
2. Preliminaries
In this section, we briefly recall necessary definitions and basic lemmas from the first part [10]. Throughout this paper, we denote by $\mathbb{Z}$ (resp. $\mathbb{N}$, $\mathbb{P}$) the set of integers (resp. nonnegative integers, positive integers). For $n\in\mathbb{N}$ we write $[n]=\{1,2,\dots,n\}$. For a statement $p$, we define $\chi(p)$ to be $1$ if $p$ is true and 00 otherwise.
A partition is a weakly decreasing sequence $\lambda=(\lambda_{1},\dots,\lambda_{\ell})$ of positive integers. We call each $\lambda_{i}$ a part of $\lambda$ and write $\ell(\lambda)=\ell$ for the number of parts. If $i>\ell(\lambda)$, we define $\lambda_{i}=0$. The set of all partitions is denoted by $\operatorname{Par}$ and the set of partitions with at most $n$ parts is denoted by $\operatorname{Par}_{n}$.
Let $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\operatorname{Par}_{n}$. We will identify the partition $\lambda$ with its Young diagram, i.e., $\lambda=\{(i,j)\in\mathbb{Z}^{2}:1\leq i\leq n,1\leq j\leq\lambda_{i}\}$. Each $(i,j)\in\lambda$ is called a cell of $\lambda$. The transpose $\lambda^{\prime}$ of $\lambda$ is the partition whose Young diagram is $\{(j,i)\in\mathbb{Z}^{2}:1\leq i\leq n,1\leq j\leq\lambda_{i}\}$. We use “the English notation” to visualize partitions; see Figure 1.
[…]
Figure 1. The Young diagram of $\lambda=(4,3,1)$ on the left and its transpose $\lambda^{\prime}=(3,2,2,1)$ on the right.
For two partitions $\lambda$ and $\mu$, we write $\mu\subseteq\lambda$ if the Young diagram of $\mu$ is contained in that of $\lambda$. In this case the skew shape ${\lambda/\mu}$ is defined to be the set-theoretic difference $\lambda-\mu$ of the Young diagrams. The size of ${\lambda/\mu}$ is the number $|{\lambda/\mu}|$ of cells in ${\lambda/\mu}$. The partition $\lambda$ will also be considered as the skew shape $\lambda/\emptyset$, where $\emptyset$ is the empty partition.
Let $\bm{x}=(x_{1},x_{2},\dots)$ be a set of formal variables. A formal power series $f(\bm{x})$ in $\bm{x}$ is called a symmetric function if it is invariant under permuting the variables. Let $\bm{x}_{n}=(x_{1},\dots,x_{n})$ and define $f(\bm{x}_{n})$ to be the formal power series $f(\bm{x})$ with substitution $x_{i}=0$ for $i>n$. In the literature, symmetric functions often mean symmetric formal power series of bounded degree, but in this paper we do not require the bounded degree condition. We refer the reader to [18, Chapter 7] and [16] for more details on symmetric functions.
The complete homogeneous symmetric function $h_{n}(\bm{x})$ and the elementary symmetric function $e_{n}(\bm{x})$ are defined by
$h_{n}(\bm{x})=\sum_{i_{1}\leq\dots\leq i_{n}}x_{i_{1}}\cdots x_{i_{n}},\qquad e_{n}(\bm{x})=\sum_{i_{1}<\dots<i_{n}}x_{i_{1}}\cdots x_{i_{n}},$
where we use the convention that $h_{0}(\bm{x})=e_{0}(\bm{x})=1$ and $h_{n}(\bm{x})=e_{n}(\bm{x})=0$ for $n<0$. For a partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell})$, we define $h_{\lambda}(\bm{x})=h_{\lambda_{1}}(\bm{x})\cdots h_{\lambda_{\ell}}(\bm{x})$ and $e_{\lambda}(\bm{x})=e_{\lambda_{1}}(\bm{x})\cdots e_{\lambda_{\ell}}(\bm{x})$.
Plethystic substitution is an important notion in symmetric function theory, which will play a key role in this paper. The plethystic substitution $f[Z]$ is defined for any symmetric function $f(\bm{x})$ and a formal power series $Z$ in any variables. However, in this paper, we only need the special cases that $f(\bm{x})$ is either $h_{n}(\bm{x})$ or $e_{n}(\bm{x})$, and $Z$ is a sum of variables with coefficients $1$ or $-1$. Hence one may take the following proposition as the definition of plethystic substitution. We refer the reader to the first part [10] for the precise definition. See also [15, 16] for more details on plethystic substitution.
Proposition 2.1.
Let $Y=y_{1}+\cdots+y_{r}$ and $Z=z_{1}+\cdots+z_{s}$, where $y_{1},\dots,y_{r},z_{1},\dots,z_{s}$ are formal variables. Then
$h_{n}[Y]$ $=h_{n}(y_{1},\dots,y_{r}),$ $e_{n}[Y]$ $=e_{n}(y_{1},\dots,y_{r}),$
$h_{n}[-Y]$ $=(-1)^{n}e_{n}[Y],$ $e_{n}[-Y]$ $=(-1)^{n}h_{n}[Y],$
$h_{n}[Y+Z]$ $=\sum_{a+b=n}h_{a}[Y]h_{b}[Z],$ $e_{n}[Y+Z]$ $=\sum_{a+b=n}e_{a}[Y]e_{b}[Z],$
$h_{n}[Y-Z]$ $=\sum_{a+b=n}h_{a}[Y]h_{b}[-Z],$ $e_{n}[Y-Z]$ $=\sum_{a+b=n}e_{a}[Y]e_{b}[-Z].$
Here, it is possible that $r=\infty$ or $s=\infty$. If $r=\infty$, then it means that $Y=y_{1}+y_{2}+\cdots$ and $h_{n}[Y]=h_{n}(y_{1},y_{2},\dots)$.
We recall the notation $\ominus$ introduced in the first part [10]:
$h_{n}[Y\ominus Z]=\sum_{a-b=n}h_{a}[Y]h_{b}[Z],\qquad e_{n}[Y\ominus Z]=\sum_{a-b=n}e_{a}[Y]e_{b}[Z].$
For integers $r,s$, and $n$, let
$X=x_{1}+x_{2}+\cdots,\qquad X_{n}=x_{1}+x_{2}+\dots+x_{n},\qquad X_{[r,s]}=x_{r}+x_{r+1}+\dots+x_{s},$
$A_{n}=\alpha_{1}+\alpha_{2}+\dots+\alpha_{n},\qquad B_{n}=\beta_{1}+\beta_{2}+\dots+\beta_{n},$
where $X_{n}=A_{n}=B_{n}=0$ if $n\leq 0$ and $X_{[r,s]}=0$ if $r>s$.
The next four basic lemmas will be used later.
Lemma 2.2 ([11, Lemma 5.2]).
Suppose that $Z$ is any formal power series, $z$ is a single variable, and $c\in\{0,1\}$. Then we have
$h_{m}[Z]=h_{m}[Z-cz]+czh_{m-1}[Z].$
Lemma 2.3 ([11, Lemma 4.7]).
Suppose that $Z$ is any formal power series, $z$ is a single variable, and $c\in\{0,1\}$. Then we have
$e_{m}[Z]=e_{m}[Z-cz]+cze_{m-1}[Z-cz]\quad\mbox{and}\quad e_{m}[Z]=e_{m}[Z+cz]-cze_{m-1}[Z].$
Lemma 2.4 ([10, Lemma 2.5]).
Let $\lambda,\mu\in\operatorname{Par}_{n}$ with $\mu\not\subseteq\lambda$ and let $Z_{i,j}$ be any formal power series for $1\leq i,j\leq n$. Then
$\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[Z_{i,j}]\right)_{i,j=1}^{n}=\det\left(e_{\lambda_{i}-\mu_{j}-i+j}[Z_{i,j}]\right)_{i,j=1}^{n}=0.$
Lemma 2.5 ([10, Lemma 2.5]).
Let $\lambda,\mu\in\operatorname{Par}_{n}$ with $\mu\subseteq\lambda$ and let $Z_{i,j}$ be any formal power series for $1\leq i,j\leq n$. If $1\leq k\leq n$ is an integer such that $\lambda_{k}=\mu_{k}$, then
$\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[Z_{i,j}]\right)_{i,j=1}^{n}=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[Z_{i,j}]\right)_{i,j=1}^{k-1}\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[Z_{i,j}]\right)_{i,j=k+1}^{n}.$
We finish this section with the following definition, which will be used frequently in this paper.
Definition 2.6.
A reverse partition is a weakly increasing sequence of positive integers $\bm{s}=(s_{1},\dots,s_{n})$. Denote by $\operatorname{RPar}_{n}$ the set of reverse partitions with $n$ elements. For convenience in our arguments, we define $s_{0}=0$ for $\bm{s}=(s_{1},\dots,s_{n})\in\operatorname{RPar}_{n}$. For $\bm{r}=(r_{1},\dots,r_{n}),\bm{s}=(s_{1},\dots,s_{n})\in\mathbb{P}^{n}$, we write $\bm{r}\leq\bm{s}$ if $r_{i}\leq s_{i}$ for all $1\leq i\leq n$.
3. Flagged Grothendieck polynomials
In this section we give a combinatorial model for the refined canonical stable Grothendieck polynomials $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ using marked multiset-valued tableaux. To this end we introduce a flagged version of $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ using marked multiset-valued tableaux and prove a Jacobi–Trudi-like formula for this flagged version, which reduces to the Jacobi–Trudi-like formula for $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ in Theorem 1.2. More generally, we consider two flagged versions of $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ and extend the partition $\lambda$ to a skew shape. In the last part of this section we show that one of our results implies a result of Matsumura [17] on a Jacobi–Trudi-like formula for a flagged Grothendieck polynomial.
To begin with, we introduce the following definition:
Definition 3.1.
A marked multiset-valued tableau of shape ${\lambda/\mu}$ is a filling $T$ of ${\lambda/\mu}$ with multisets such that
  • $T(i,j)$ is a nonempty (finite) multiset $\{a_{1}\leq\dots\leq a_{k}\}$ of positive integers, in which each integer $a_{i}$ may be marked if $i\geq 2$ and $a_{i-1}<a_{i}$, and
  • $\max(T(i,j))\leq\min(T(i,j+1))$ and $\max(T(i,j))<\min(T(i+1,j))$ if $(i,j),(i,j+1)\in{\lambda/\mu}$ and $(i,j),(i+1,j)\in{\lambda/\mu}$, respectively.
Let $\operatorname{MMSVT}({\lambda/\mu})$ denote the set of marked multiset-valued tableaux of shape ${\lambda/\mu}$.
For $T\in\operatorname{MMSVT}({\lambda/\mu})$, let $\bm{x}^{T(i,j)}=x_{1}^{m_{1}}x_{2}^{m_{2}}\cdots$, where $m_{k}$ is the total number of appearances of (unmarked integers) $k$ and (marked integers) $k^{*}$ in $T(i,j)$, and let $\operatorname{unmarked}(T(i,j))$ (resp. $\operatorname{marked}(T(i,j))$) denote the number of unmarked (resp. marked) integers in $T(i,j)$. We define
$\operatorname{wt}(T)=\prod_{(i,j)\in{\lambda/\mu}}\bm{x}^{T(i,j)}\alpha_{j}^{\operatorname{unmarked}(T(i,j))-1}(-\beta_{i})^{\operatorname{marked}(T(i,j))}.$
For example, see Figure 2.
[…]
1 3,3
Figure 2. An example of $T\in\operatorname{MMSVT}((3,2)/(1))$, where the marked integers are indicated with *. The weight of $T$ is given by $\operatorname{wt}(T)=x_{1}^{2}x_{2}^{4}x_{3}^{2}x_{4}\alpha_{2}^{2}\alpha_{3}(-\beta_{1})^{2}$.
The main goal of this section is to prove the following combinatorial interpretation for $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$.
Theorem 3.2.
For a partition $\lambda$, we have
$G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})=\sum_{T\in\operatorname{MMSVT}(\lambda)}\operatorname{wt}(T).$
Remark 3.3.
[…]
Remark 3.4.
We note that the crystal operators introduced by Hawkes and Scrimshaw [9, Theorem 4.6] on MMSVT (when translated from hook-valued tableaux under the natural bijection) preserve the statistics that determine $\bm{\alpha}$ and $\bm{\beta}$, which gives an alternative proof of the Schur positivity of $G_{\lambda}(\bm{x};\bm{\alpha},-\bm{\beta})$ (private communication with Travis Scrimshaw).
Our strategy for proving Theorem 3.2 is to introduce a flagged version which allows us to use induction. We note that a row-flagged version with partition shape is sufficient to prove this theorem, but for completeness we also consider a column-flagged version and we generalize the partition shape to any skew shape.
Definition 3.5.
Let $\bm{r}=(r_{1},\dots,r_{n})\in\mathbb{N}^{n}$ and $\bm{s}=(s_{1},\dots,s_{n})\in\mathbb{N}^{n}$. Let $\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})$ denote the set of $T\in\operatorname{MMSVT}({\lambda/\mu})$ such that $r_{i}\leq\min(T(i,j))$ and $\max(T(i,j))\leq s_{i}$ for all $(i,j)\in{\lambda/\mu}$. Similarly, $\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s})}({\lambda/\mu})$ denotes the set of $T\in\operatorname{MMSVT}({\lambda/\mu})$ such that $r_{j}\leq\min(T(i,j))$ and $\max(T(i,j))\leq s_{j}$ for all $(i,j)\in{\lambda/\mu}$. We define the row-flagged and column-flagged refined canonical stable Grothendieck polynomials by
$G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\sum_{T\in\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})}\operatorname{wt}(T),$
$G_{\lambda/\mu}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\sum_{T\in\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s})}({\lambda/\mu})}\operatorname{wt}(T).$
Remark 3.6.
Recall that if $T\in\operatorname{MMSVT}({\lambda/\mu})$, then every cell of $T$ consists only of positive integers. Therefore, in Definition 3.5, it suffices to consider $\bm{r},\bm{s}\in\mathbb{P}^{n}$. However, we extend the definition to $\bm{r},\bm{s}\in\mathbb{N}^{n}$ because later in this section, for example in Lemma 3.15, we will prove recurrences for $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ in which an element $s_{k}$ of $\bm{s}$ is decreased by $1$.
To illustrate, suppose $s_{k}=0$. If $\mu_{k}<\lambda_{k}$, then there is a cell in row $k$, say $(i,k)$, in $\lambda/\mu$. Thus, if $T\in\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})$, then every integer in $T(i,k)$ must be at most $s_{k}=0$. However, since $T(i,k)$ must consist of positive integers, there is no such $T$, hence $\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})=\emptyset$. On the other hand, if $\mu_{k}=\lambda_{k}$, then there are no cells in row $k$ of $\lambda/\mu$. Therefore the upper bound $s_{k}=0$ is irrelevant and we may have $\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})\neq\emptyset$.
Now we state Jacobi–Trudi-like formulas for $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $G_{\lambda/\mu}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ whose proofs are given in Sections 3.1 and 3.2 respectively. Here we use the convention that if $r>s$, then the product $\prod_{i=r}^{s}a_{i}$ is defined to be $1$.
Theorem 3.7.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. If $r_{i}\leq r_{i+1}$ and $s_{i}\leq s_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$ for $11\leq i\leq n-$ 1, then
(3.1) $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=C\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j})]\right)_{i,j=1}^{n},$
where
$C=\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l}).$
Theorem 3.8.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. If $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}$ and $s_{i}-\lambda_{i}\leq s_{i+1}-\lambda_{i+1}+1$ whenever $\mu_{i}<\lambda_{i+1}$ for $11\leq i\leq n-$ 1, then
(3.2) $G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=D\det\left(e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{i-1}-A_{j}-B_{\lambda_{i}}+B_{\mu_{j}})]\right)_{i,j=1}^{n},$
where
$D=\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\alpha_{i}x_{l})^{-1}.$
Remark 3.9.
The condition $\mu\subseteq\lambda$ in Theorems 3.7 and 3.8 is necessary. Indeed, if $\lambda=(1),\mu=(2),\bm{r}=(1),\bm{s}=(1)$, then the left hand side of (3.1) (resp. (3.2)) is zero, while the right hand side is $\beta_{1}-\alpha_{2}$ (resp. $\beta_{2}-\alpha_{1}$).
By specializing Theorem 3.7 we obtain the following corollary, which implies the combinatorial model for $G_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ in Theorem 3.2 as $n\to\infty$.
Corollary 3.10.
For a partition $\lambda\in\operatorname{Par}_{n}$, we have
$G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})=\sum_{T\in\operatorname{MMSVT}(\lambda),\max(T)\leq n}\operatorname{wt}(T),$
where $\max(T)$ is the largest integer appearing in $T$.
Proof.
Let $\bm{r}=(1^{n})$ and $\bm{s}=(n^{n})$, where $(a^{n})$ means the sequence $(a,a,\dots,a)$ with $n$ $a$’s. By definition, we have
$G_{\lambda}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=\sum_{T\in\operatorname{MMSVT}(\lambda),\max(T)\leq n}\operatorname{wt}(T).$
On the other hand, by Theorems 3.7 and 1.2, we have
$G_{\lambda}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\prod_{i,j=1}^{n}(1-\beta_{i}x_{j})\det\left(h_{\lambda_{i}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{\lambda_{i}}-B_{i-1}+B_{j})]\right)_{i,j=1}^{n}$
$=G_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta}).$
Combining the above two equations gives the corollary. ∎
Remark 3.11.
Note that the Jacobi–Trudi formula for the Schur function $s_{{\lambda/\mu}}(\bm{x})$ of shape ${\lambda/\mu}$ does not depend on the choice of $n$ for $n>\ell(\lambda)$, that is,
$s_{\lambda/\mu}(\bm{x})=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}(\bm{x})\right)_{i,j=1}^{n}=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}(\bm{x})\right)_{i,j=1}^{\ell(\lambda)}.$
This can be easily proved using the fact that $h_{m}(\bm{x})=0$ for $m<0$. However, since $h_{m}[X\ominus Y]\not=0$ for $m<0$ in general, it is not obvious that the Jacobi–Trudi-like formula for $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ in Theorem 3.7 does not depend on the choice of $n$ for $n>\ell(\lambda)$. It is still possible to prove the independence of $n$ directly by showing that for $i$ and $j$ satisfying $j<i$ and $\lambda_{i}=\mu_{i}=0$, the $(i,j)$-entry of the matrix in (3.1) is zero.
In the next two subsections we prove Theorems 3.7 and 3.8.
3.1. Proof of Theorem 3.7: row-flagged Grothendieck polynomials
For $\lambda,\mu\in\operatorname{Par}_{n},$ and $\bm{r},\bm{s}\in\mathbb{N}^{n}$, we define
(3.3) $\tilde{G}_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=C\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j})]\right)_{i,j=1}^{n},$
where $C=\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l})$. We prove the Jacobi–Trudi-like identity (3.1) in Theorem 3.7 by showing that $G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $\tilde{G}_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ satisfy the same recurrence relations and initial conditions. The idea is based on a proof of the Jacobi–Trudi-like formula for flagged Schur functions due to Wachs [19].
Throughout this section we assume that $n$ is fixed and $\epsilon_{k}$ denotes the 0-1 sequence of length $n$ whose $k$th entry is $1$ and all other entries are 00.
In order to prove Theorem 3.7 we need several lemmas. Let $\phi_{\bm{\beta}}$ be the substitution homomorphism that replaces $\beta_{i}$ by $\beta_{i+1}$ for each $i\geq 1$. For example, $\phi_{\bm{\beta}}(x_{1}\beta_{1}\beta_{3}+x_{2}\alpha_{2}\beta_{1}+\alpha_{3})=x_{1}\beta_{2}\beta_{4}+x_{2}\alpha_{2}\beta_{2}+\alpha_{3}$.
Lemma 3.12.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{N}^{n}$ with $\mu\subseteq\lambda$ and let $11\leq k\leq n-$ 1 be an integer such that $\mu_{k}\geq\lambda_{k+1}$. Then
$G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=G_{{\lambda}^{(1)}/{\mu}^{(1)}}^{\operatorname{row}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\cdot\phi_{\bm{\beta}}^{k}\left(G_{{\lambda}^{(2)}/{\mu}^{(2)}}^{\operatorname{row}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right),$
$\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\tilde{G}_{{\lambda}^{(1)}/{\mu}^{(1)}}^{\operatorname{row}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\cdot\phi_{\bm{\beta}}^{k}\left(\tilde{G}_{{\lambda}^{(2)}/{\mu}^{(2)}}^{\operatorname{row}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right),$
where for a sequence $\bm{t}=(t_{1},\ldots,t_{n})$ we denote $\bm{t}^{(1)}=(t_{1},\ldots,t_{k})$ and $\bm{t}^{(2)}=(t_{k+1},\ldots,t_{n})$.
Proof.
Observe that ${\lambda/\mu}$ is the union of ${\lambda}^{(1)}/{\mu}^{(1)}$ and ${\lambda}^{(2)}/{\mu}^{(2)}$ whose column indices are disjoint. Thus, the first identity directly follows from the definition of $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$.
For the second identity, it is easy to check that both sides have the same factor $\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l}).$ We will show that the matrix in (3.3) is a block upper triangular matrix with diagonal blocks of sizes $k\times k$ and $(n-k)\times(n-k)$. Since $\mu_{1}\geq\cdots\geq\mu_{k}\geq\lambda_{k+1}\geq\cdots\geq\lambda_{n}$, we have $\lambda_{i}-\mu_{j}-i+j<0$ for all $k+1\leq i\leq n$ and $1\leq j\leq k$. Furthermore, for such $i$ and $j$, the $(i,j)$-entry of the matrix in (3.3) is equal to
$\sum_{\ell\geq 0}h_{\lambda_{i}-\mu_{j}-i+j+\ell}\left[X_{[r_{j},s_{i}]}\right]h_{\ell}\left[A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j}\right]$
$=\sum_{\ell\geq-\lambda_{i}+\mu_{j}+i-j}h_{\lambda_{i}-\mu_{j}-i+j+\ell}\left[X_{[r_{j},s_{i}]}\right]h_{\ell}\left[A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j}\right]$
$=\sum_{\ell\geq-\lambda_{i}+\mu_{j}+i-j}h_{\lambda_{i}-\mu_{j}-i+j+\ell}\left[X_{[r_{j},s_{i}]}\right](-1)^{\ell}e_{\ell}\left(\alpha_{\lambda_{i}+1},\dots,\alpha_{\mu_{j}},\beta_{j+1},\dots,\beta_{i-1}\right),$
which is equal to 00 because for each $\ell\geq-\lambda_{i}+\mu_{j}+i-j$, the number of variables in $e_{\ell}(\alpha_{\lambda_{i}+1},\dots,\alpha_{\mu_{j}},\beta_{j+1},\dots,\beta_{i-1})$ is $(\mu_{j}-\lambda_{i})+(i-1-j)<\ell$. Thus the matrix in (3.3) is a block upper triangular matrix and we obtain the second identity. ∎
The following two lemmas show that $G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $\tilde{G}_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ have the same initial conditions.
Lemma 3.13.
For $\lambda\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{N}^{n}$,
$G_{\lambda/\lambda}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{\lambda/\lambda}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=1.$
Proof.
Since $\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}(\lambda/\lambda)$ has only one element, namely the empty tableau, we have $G_{\lambda/\lambda}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=1$.
For the second equality, let $H_{i,j}=h_{\lambda_{i}-\lambda_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{\lambda_{i}}-A_{\lambda_{j}}-B_{i-1}+B_{j})]$ so that $\tilde{G}_{\lambda/\lambda}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=C\det(H_{i,j})_{i,j=1}^{n}$, where $C=\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l})$. If we set $\mu=\lambda$, then since $\mu_{k}=\lambda_{k}\geq\lambda_{k+1}$ for all $11\leq k\leq n-$ 1, we have $C\det(H_{i,j})_{i,j=1}^{n}=C\prod_{i=1}^{n}H_{i,i}$ by iterated application of Lemma 3.12. Since
(3.4) $H_{i,i}=\sum_{\ell\geq 0}h_{\ell}[X_{[r_{i},s_{i}]}]h_{\ell}[\beta_{i}]=\sum_{\ell\geq 0}h_{\ell}[X_{[r_{i},s_{i}]}]\beta_{i}^{\ell}=\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l})^{-1},$
we obtain $\tilde{G}_{\lambda/\lambda}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=C\det(H_{i,j})_{i,j=1}^{n}=C\cdot C^{-1}=1$. Note that if $r_{i}>s_{i}$, then $X_{[r_{i},s_{i}]}=0$, and therefore (3.4) is still true since $\sum_{\ell\geq 0}h_{\ell}[X_{[r_{i},s_{i}]}]\beta_{i}^{\ell}=1=\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l})^{-1}$. ∎
Lemma 3.14.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\operatorname{RPar}_{n}$ with $\mu\subseteq\lambda$. If $s_{k}<r_{k}$ and $\mu_{k}<\lambda_{k}$ for some $1\leq k\leq n$, then
$G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=0.$
Proof.
Since $\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}(\lambda/\mu)=\emptyset$, we have $G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=0$. To prove the second equality, we claim that for all $1\leq i\leq k$ and $k\leq j\leq n$, the $(i,j)$-entry of the matrix in (3.3) is $0,$ that is,
$\sum_{\ell\geq 0}h_{\lambda_{i}-\mu_{j}-i+j+\ell}\left[X_{[r_{j},s_{i}]}\right]h_{\ell}\left[A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j}\right]=0.$
Observe that $s_{i}\leq s_{k}<r_{k}\leq r_{j}$ and $\mu_{j}\leq\mu_{k}<\lambda_{k}\leq\lambda_{i}$ by the assumptions of the lemma. Thus $h_{\lambda_{i}-\mu_{j}-i+j+\ell}[X_{[r_{j},s_{i}]}]=0$ for $\ell\geq 0$ since $X_{[r_{j},s_{i}]}=0$ and $\lambda_{i}-\mu_{j}-i+j+\ell>0$. Therefore, the claim is true, which implies $\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=0$ (because if an $n\times n$-matrix $(a_{i,j})$ satisfies $a_{i,j}=0$ for all $1\leq i\leq k$ and $k\leq j\leq n$, then its determinant is 0). ∎
Recall that for $\lambda\in\operatorname{Par}_{n}$ and $\bm{s}\in\operatorname{RPar}_{n}$ we assume $\lambda_{n+1}=0$ and $s_{0}=0$. The following two lemmas show that $G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $\tilde{G}_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ have the same recursions.
Lemma 3.15.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\operatorname{RPar}_{n}$ with $\mu\subseteq\lambda$. Let $1\leq k\leq n$ be an integer satisfying the following conditions:
  1. (1)
    $\mu_{k}<\lambda_{k}$,
  2. (2)
    $r_{k}\leq s_{k}$,
  3. (3)
    $s_{k-1}<s_{k}$, and
  4. (4)
    $\lambda_{k}>\lambda_{k+1}$.
Then
$G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\frac{1-x_{s_{k}}\beta_{k}}{1-x_{s_{k}}\alpha_{\lambda_{k}}}G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})+\frac{x_{s_{k}}}{1-x_{s_{k}}\alpha_{\lambda_{k}}}G_{(\lambda-\epsilon_{k})/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta}),$
$\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\frac{1-x_{s_{k}}\beta_{k}}{1-x_{s_{k}}\alpha_{\lambda_{k}}}\tilde{G}_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})+\frac{x_{s_{k}}}{1-x_{s_{k}}\alpha_{\lambda_{k}}}\tilde{G}_{(\lambda-\epsilon_{k})/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Moreover, if $s_{k}=1$, then
$G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0.$
Proof.
Considering the elements in the cell $(k,\lambda_{k})$, we partition $\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})$ into two sets:
$S_{1}$ =$=\left\{T\in\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu}):T(k,\lambda_{k})\text{ contains an integer }\ell<s_{k}\right\},$
[…]
For each $T\in S_{1}$, by extracting all $s_{k}$’s from $T(k,\lambda_{k})$, we obtain
$\sum_{T\in S_{1}}\operatorname{wt}(T)=\frac{1-x_{s_{k}}\beta_{k}}{1-x_{s_{k}}\alpha_{\lambda_{k}}}G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Note that if $s_{k}=1$, then $S_{1}=\emptyset$ and, by Remark 3.6, $\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}({\lambda/\mu})=\emptyset$. Hence in this case both sides of the above identity are equal to 00 and in particular we have $G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0$. For each $T\in S_{2}$, extracting all $s_{k}$’s from $T(k,\lambda_{k})$ and deleting the cell $(k,\lambda_{k})$, we obtain
$\sum_{T\in S_{2}}\operatorname{wt}(T)=\frac{x_{s_{k}}}{1-x_{s_{k}}\alpha_{\lambda_{k}}}G_{(\lambda-\epsilon_{k})/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Adding the above two equations gives the first identity.
For the second identity let
$H_{i,j}$ $=h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j})],$
$F_{i,j}$ $=h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}-\chi(i=k)]}\ominus(A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j})],$
$G_{i,j}$ $=h_{\lambda_{i}-\mu_{j}-i+j-\chi(i=k)}[X_{[r_{j},s_{i}]}\ominus(A_{\lambda_{i}-\chi(i=k)}-A_{\mu_{j}}-B_{i-1}+B_{j})].$
By (3.3) the second identity is written as
$C\det(H_{i,j})=\frac{C}{1-x_{s_{k}}\alpha_{\lambda_{k}}}\det(F_{i,j})+\frac{Cx_{s_{k}}}{1-x_{s_{k}}\alpha_{\lambda_{k}}}\det(G_{i,j}),$
where $C=\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l})$. Since $H_{i,j}=F_{i,j}=G_{i,j}$ for all $1\leq i,j\leq n$ with $i\neq k$, by the linearity of a determinant in its $k$th row, it suffices to show that
(3.5) $H_{k,j}=\frac{1}{1-x_{s_{k}}\alpha_{\lambda_{k}}}F_{k,j}+\frac{x_{s_{k}}}{1-x_{s_{k}}\alpha_{\lambda_{k}}}G_{k,j}$
for all $1\leq j\leq n$. For brevity let $m=\lambda_{k}-\mu_{j}-k+j$ and $Z=A_{\lambda_{k}}-A_{\mu_{j}}-B_{k-1}+B_{j}$ so that
(3.6) $H_{k,j}$ $=\sum_{\ell\geq 0}h_{m+\ell}\left[X_{[r_{j},s_{k}]}\right]h_{\ell}[Z],$
(3.7) $F_{k,j}$ $=\sum_{\ell\geq 0}h_{m+\ell}\left[X_{[r_{j},s_{k}-1]}\right]h_{\ell}[Z],$
(3.8) $G_{k,j}$ $=\sum_{\ell\geq 0}h_{m+\ell-1}\left[X_{[r_{j},s_{k}]}\right]h_{\ell}[Z-\alpha_{\lambda_{k}}].$
We consider the two cases $r_{j}\leq s_{k}$ and $r_{j}>s_{k}$. First, suppose $r_{j}\leq s_{k}$. We will express $F_{k,j}$ in terms of $H_{k,j}$ and $G_{k,j}$. Since $h_{m+\ell}[X_{[r_{j},s_{k}-1]}]=h_{m+\ell}[X_{[r_{j},s_{k}]}-x_{s_{k}}]$, by Lemma 2.2 we have
$h_{m+\ell}\left[X_{[r_{j},s_{k}-1]}\right]$ $=h_{m+\ell}\left[X_{[r_{j},s_{k}]}\right]-x_{s_{k}}h_{m+\ell-1}\left[X_{[r_{j},s_{k}]}\right],$
$h_{\ell}[Z]$ $=h_{\ell}[Z-\alpha_{\lambda_{k}}]+\alpha_{\lambda_{k}}h_{\ell-1}[Z].$
Thus
$F_{k,j}$ $=\sum_{\ell\geq 0}h_{m+\ell}\left[X_{[r_{j},s_{k}]}\right]h_{\ell}[Z]-x_{s_{k}}\sum_{\ell\geq 0}h_{m+\ell-1}\left[X_{[r_{j},s_{k}]}\right]h_{\ell}[Z]$
$=H_{k,j}-x_{s_{k}}\sum_{\ell\geq 0}h_{m+\ell-1}\left[X_{[r_{j},s_{k}]}\right]\left(h_{\ell}[Z-\alpha_{\lambda_{k}}]+\alpha_{\lambda_{k}}h_{\ell-1}[Z]\right)$
$=H_{k,j}-x_{s_{k}}G_{k,j}-x_{s_{k}}\alpha_{\lambda_{k}}H_{k,j},$
which is equivalent to (3.5).
Now suppose $r_{j}>s_{k}$. We will show that $H_{k,j}=F_{k,j}=G_{k,j}=0$, which implies (3.5). First observe that, since $X_{[r_{j},s_{k}]}=X_{[r_{j},s_{k}-1]}=0$ and $h_{t}[0]=\chi(t=0)$, we obtain from (3.6), (3.7), and (3.8) that
(3.9) $H_{k,j}=F_{k,j}=h_{-m}[Z],\qquad G_{k,j}=h_{1-m}[Z-\alpha_{\lambda_{k}}],$
where $m=\lambda_{k}-\mu_{j}-k+j$ and $Z=A_{\lambda_{k}}-A_{\mu_{j}}-B_{k-1}+B_{j}$ as before and $h_{t}(\bm{x})=0$ for $t<0$. Since $r_{k}\leq s_{k}<r_{j}$ and $\bm{r}\in\operatorname{RPar}_{n}$, we must have $j>k$. Thus $\lambda_{k}>\mu_{k}\geq\mu_{j}$ and hence $m=(\lambda_{k}-\mu_{j})+(j-k)\geq 1+1$, which together with (3.9) implies $H_{k,j}=F_{k,j}=G_{k,j}=0$. Thus, (3.5) holds.
To complete the proof, it remains to show that $\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0$ for the case $s_{k}=1$. If $s_{k}=1$, then $r_{j}>s_{k}-1=0$. Applying the same argument as before, we have $F_{k,j}=0$ for any $j$ and hence $\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0.$
Lemma 3.16.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\operatorname{RPar}_{n}$ with $\mu\subseteq\lambda$. Let $11\leq k\leq n-$ 1 be an integer satisfying the following conditions:
  1. (1)
    $\mu_{k}<\lambda_{k}$,
  2. (2)
    $r_{k}\leq s_{k}$,
  3. (3)
    $s_{k}=s_{k+1}$, and
  4. (4)
    $\lambda_{k}=\lambda_{k+1}$.
Then
$G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta}),$
$\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Moreover, if $s_{k}=1$, then
$G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0.$
Proof.
Let $T\in\operatorname{MMSVT}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})$. Then
$\max(T(k,\lambda_{k}))<\min(T(k+1,\lambda_{k}))=\min(T(k+1,\lambda_{k+1}))\leq s_{k+1}=s_{k}.$
Thus every integer in the $k$th row of $T$ is at most $s_{k}-1$, which shows the first identity. As before, we have $G_{{\lambda/\mu}}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0$ if $s_{k}=1$ by Remark 3.6.
For the second identity let
$H_{i,j}$ $=h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j})],$
$F_{i,j}$ $=h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}-\chi(i=k)]}\ominus(A_{\lambda_{i}}-A_{\mu_{j}}-B_{i-1}+B_{j})].$
Dividing both sides by $\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\beta_{i}x_{l})$, the second identity can be rewritten as
(3.10) $\det(H_{i,j})=\frac{1}{1-x_{s_{k}}\beta_{k}}\det(F_{i,j}).$
We claim that for all $1\leq j\leq n$,
(3.11) $F_{k,j}=(1-x_{s_{k}}\beta_{k})H_{k,j}-x_{s_{k}}H_{k+1,j}.$
Since $F_{i,j}=H_{i,j}$ for all $1\leq i,j\leq n$ with $i\neq k$, by the linearity of a determinant in its $k$th row, the claim implies (3.10). To prove the claim, we compare $F_{k,j}$ and $H_{k,j}$. Let $m=\lambda_{k}-\mu_{j}-k+j$ and $Z=A_{\lambda_{k}}-A_{\mu_{j}}-B_{k-1}+B_{j}$ for brevity. We consider the two cases $r_{j}\leq s_{k}$ and $r_{j}>s_{k}$.
First, suppose $r_{j}\leq s_{k}$. By Lemma 2.2, we have
$h_{m+\ell}\left[X_{[r_{j},s_{k}-1]}\right]$ $=h_{m+\ell}\left[X_{[r_{j},s_{k}]}\right]-x_{s_{k}}h_{m+\ell-1}\left[X_{[r_{j},s_{k}]}\right],$
$h_{\ell}[Z]$ $=h_{\ell}[Z-\beta_{k}]+\beta_{k}h_{\ell-1}[Z].$
The claim then follows from similar computations as in the proof of Lemma 3.15 with the above two equations since $s_{k}=s_{k+1}$ and $\lambda_{k}=\lambda_{k+1}$ easily lead to $H_{k+1,j}=\sum_{\ell\geq 0}h_{m+\ell-1}[X_{[r_{j},s_{k}]}]h_{\ell}[Z-\beta_{k}]$.
Now suppose $r_{j}>s_{k}$. By the same argument as in the proof of Lemma 3.15, we obtain $H_{k,j}=F_{k,j}=0$ and also $H_{k+1,j}=0$. This shows the claim (3.11).
To complete the proof, it remains to show that $\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0$ for the case $s_{k}=1$. In this case, since $r_{j}>s_{k}-1=0$, one can apply the same argument above to obtain $F_{k,j}=0$ for any $j$, which gives $\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0$. ∎
We are now ready to prove Theorem 3.7, which is restated as the following theorem.
Theorem 3.17.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. If $r_{i}\leq r_{i+1}$ and $s_{i}\leq s_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$ for $11\leq i\leq n-$ 1, then
$G_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Proof.
We proceed by strong induction on $N:=n+|{\lambda/\mu}|+\sum_{i=1}^{n}s_{i}\geq 1$, where it is unnecessary to check the base case. Suppose that the theorem holds for all integers less than $N$. If $\mu=\lambda$, then it follows from Lemma 3.13. Therefore we may assume $\mu\neq\lambda$. We consider the following two cases.
Case 1: Suppose that there is an integer $11\leq t\leq n-$ 1 satisfying $\mu_{t}\geq\lambda_{t+1}$. We set $\gamma^{(1)}=(\gamma_{1},\dots,\gamma_{t})$ and $\gamma^{(2)}=(\gamma_{t+1},\dots,\gamma_{n})$ for each $\gamma\in\{\lambda,\mu,\bm{r},\bm{s}\}$. By Lemma 3.12, the statement is separated into the cases $(\lambda^{(i)},\mu^{(i)},\bm{r}^{(i)},\bm{s}^{(i)})$ for $i=1,2$. Each case is then covered by the induction hypothesis.
Case 2: Suppose that $\mu_{t}<\lambda_{t+1}$ for all $11\leq t\leq n-$ 1 (or $n=1$). Then by the assumption in the theorem we have $\bm{r},\bm{s}\in\operatorname{RPar}_{n}$. Moreover, since $\lambda\in\operatorname{Par}_{n}$, we have $\mu_{i}<\lambda_{i}$ for all $1\leq i\leq n$ (if $n=1$, this follows from the assumptions $\mu\subseteq\lambda$ and $\mu\neq\lambda$). Since $\bm{s}\in\operatorname{RPar}_{n}$ (with $s_{0}=0$), we can find a unique integer $1\leq k\leq n$ satisfying $s_{k-1}<s_{k}=\dots=s_{n}$. If $s_{k}<r_{k}$, then the statement follows from Lemma 3.14. Thus we now assume $s_{k}\geq r_{k}$.
If $\lambda_{k}>\lambda_{k+1}$, then we apply Lemma 3.15. If $\lambda_{k}=\lambda_{k+1}$, then $k<n$ because otherwise we would have $00\leq\mu_{n}<\lambda_{n}=\lambda_{n+1}=$ 0, which is a contradiction. Therefore we can apply Lemma 3.16. In either case the statement reduces to some cases with a smaller $N$, and therefore the induction hypothesis applies.
The above two cases show that the statement holds for $N$, and the theorem follows by induction. ∎
3.2. Proof of Theorem 3.8: column-flagged Grothendieck polynomials
In this subsection we prove the Jacobi–Trudi-like identity for $G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ in Theorem 3.8. For $\lambda,\mu\in\operatorname{Par}_{n}$ and $\bm{r},\bm{s}\in\mathbb{N}^{n}$, let
(3.12) $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=D\det\left(e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{i-1}-A_{j}-B_{\lambda_{i}}+B_{\mu_{j}})]\right)_{i,j=1}^{n},$
where $D=\prod_{i=1}^{n}\prod_{l=r_{i}}^{s_{i}}(1-\alpha_{i}x_{l})^{-1}$. Then the identity in Theorem 3.8 is written as
$G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Since we prove this identity in an analogous way as in the previous subsection, most of the proofs in this subsection will be omitted and some additional explanations will be given when necessary.
As before we need several lemmas.
Lemma 3.18.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{N}^{n}$ with $\mu\subseteq\lambda$ and suppose $11\leq k\leq n-$ 1 is an integer such that $\mu_{k}\geq\lambda_{k+1}$. If $\phi_{\bm{\alpha}}$ is the map defined by shifting $\alpha_{i}$ to $\alpha_{i+1}$, then
$G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=G_{({\lambda}^{(1)})^{\prime}/({\mu}^{(1)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\cdot\phi_{\bm{\alpha}}^{k}\left(G_{({\lambda}^{(2)})^{\prime}/({\mu}^{(2)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right),$
$\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\tilde{G}_{({\lambda}^{(1)})^{\prime}/({\mu}^{(1)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\cdot\phi_{\bm{\alpha}}^{k}\left(\tilde{G}_{({\lambda}^{(2)})^{\prime}/({\mu}^{(2)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right),$
where for a sequence $\bm{t}=(t_{1},\ldots,t_{n})$ we denote $\bm{t}^{(1)}=(t_{1},\ldots,t_{k})$ and $\bm{t}^{(2)}=(t_{k+1},\ldots,t_{n})$.
Proof.
It is similar to the proof of Lemma 3.12. ∎
The following two lemmas show that $G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ have the same initial conditions.
Lemma 3.19.
For $\lambda\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{N}^{n}$,
$G_{\lambda^{\prime}/\lambda^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{\lambda^{\prime}/\lambda^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=1.$
Proof.
It is similar to the proof of Lemma 3.13. ∎
Lemma 3.20.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. Suppose $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1},s_{i}-\lambda_{i}\leq s_{i+1}-\lambda_{i+1}+1$ and $\mu_{i}<\lambda_{i+1}$ for any $11\leq i\leq n-$ 1. If $s_{k}-\lambda_{k}+1<r_{k}-\mu_{k}$ and $\mu_{k}<\lambda_{k}$ for some $1\leq k\leq n$, then
$G_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=0.$
Proof.
We first show that $G^{\operatorname{col}(\bm{r},\bm{s})}_{{\lambda^{\prime}/\mu^{\prime}}}(\bm{x};\bm{\alpha},\bm{\beta})=0$. To see this, suppose $T\in\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s})}({\lambda^{\prime}/\mu^{\prime}})$. The $k$th column of $T$ consists of $\lambda_{k}-\mu_{k}$ cells, so at least $\lambda_{k}-\mu_{k}$ distinct integers appear in the column. But this is impossible since $s_{k}-\lambda_{k}+1<r_{k}-\mu_{k}$. Thus $\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s})}({\lambda^{\prime}/\mu^{\prime}})=\emptyset$ and $G_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=0$.
For the second equality, we claim that for all $1\leq i\leq k$ and $k\leq j\leq n$, the $(i,j)$-entry of the matrix in (3.12) is $0,$ that is,
(3.13) $\sum_{\ell\geq 0}e_{\lambda_{i}-\mu_{j}-i+j+\ell}\left[X_{[r_{j},s_{i}]}\right]e_{\ell}\left[A_{i-1}-A_{j}-B_{\lambda_{i}}+B_{\mu_{j}}\right]=0.$
Observe that $\lambda_{i}-\mu_{j}-i+j\geq\lambda_{k}-\mu_{k}>0$. By the assumption $\mu_{m}<\lambda_{m+1}$ for all $m<n$ (combined with the assumption of the lemma), we have $r_{k}-\mu_{k}\leq r_{j}-\mu_{j}$ and $s_{i}-\lambda_{i}\leq s_{k}-\lambda_{k}+(k-i)$. Thus
$s_{i}-\lambda_{i}\leq s_{k}-\lambda_{k}+k-i<r_{k}-\mu_{k}+k-i-1\leq r_{j}-\mu_{j}+k-i-1\leq r_{j}-\mu_{j}+j-i-1,$
and we obtain $s_{i}-r_{j}+1<\lambda_{i}-\mu_{j}-i+j$. This implies $e_{\lambda_{i}-\mu_{j}-i+j+\ell}[X_{[r_{j},s_{i}]}]=0$ for all $\ell\geq 0$, because the number of variables is $\max(0,s_{i}-r_{j}+1)$, which is smaller than $\lambda_{i}-\mu_{j}-i+j+\ell$. Therefore the claim (3.13) holds and we obtain $\tilde{G}_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=0$. ∎
Recall that for $\lambda\in\operatorname{Par}_{n}$ and $\bm{s}\in\operatorname{RPar}_{n}$ we assume $\lambda_{n+1}=0$ and $s_{0}=0$. The following two lemmas show that $G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ and $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ have the same recursions.
Lemma 3.21.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$. Suppose $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}$ for all $11\leq i\leq n-$ 1. Let $1\leq k\leq n$ be an integer satisfying the following conditions:
  1. (1)
    $\mu_{k}<\lambda_{k}$,
  2. (2)
    $r_{k}-\mu_{k}\leq s_{k}-\lambda_{k}+1$,
  3. (3)
    $s_{k-1}-\lambda_{k-1}\leq s_{k}-\lambda_{k}$ if $k\geq 2$, and
  4. (4)
    $\lambda_{k}>\lambda_{k+1}$.
Then
$G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\frac{1-x_{s_{k}}\beta_{\lambda_{k}}}{1-x_{s_{k}}\alpha_{k}}G_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})+\frac{x_{s_{k}}}{1-x_{s_{k}}\alpha_{k}}G_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta}),$
$\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\frac{1-x_{s_{k}}\beta_{\lambda_{k}}}{1-x_{s_{k}}\alpha_{k}}\tilde{G}_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})+\frac{x_{s_{k}}}{1-x_{s_{k}}\alpha_{k}}\tilde{G}_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Moreover, if $s_{k}=1$, then
(3.14) $G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=0,$
(3.15) $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=0,$
(3.16) $G_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})$ =$=\begin{cases}0&\mbox{if $\lambda_{k}>\mu_{k}+1$},\\ G_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})&\mbox{if $\lambda_{k}=\mu_{k}+1$,}\end{cases}$
(3.17) $\tilde{G}_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})$ =$=\begin{cases}0&\mbox{if $\lambda_{k}>\mu_{k}+1$},\\ \tilde{G}_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})&\mbox{if $\lambda_{k}=\mu_{k}+1$.}\end{cases}$
Proof.
One can prove the first identity with the same argument as in the proof of Lemma 3.15. The only thing that needs a careful thought is whether adding any number of integers $s_{k}$ in the cell $(\lambda_{k},k)$ of a tableau in $\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\lambda^{\prime}/\mu^{\prime})$ or $\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}((\lambda-\epsilon_{k})^{\prime}/\mu^{\prime})$ always gives a tableau in $\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s})}(\lambda^{\prime}/\mu^{\prime})$. This is indeed true because $s_{k}\geq r_{k}$ (which follows from $\mu_{k}<\lambda_{k}$ and $r_{k}-\mu_{k}\leq s_{k}-\lambda_{k}+1$) and the maximum possible entry of $(\lambda_{k},k-1)$ in a tableau of $\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s})}(\lambda^{\prime}/\mu^{\prime})$ is $s_{k-1}-\lambda_{k-1}+\lambda_{k}\leq s_{k}$ (because of the assumption $s_{k-1}-\lambda_{k-1}\leq s_{k}-\lambda_{k}$). Note that if $k=1$, we do not have to consider the cell $(\lambda_{k},k-1)$ because $(\lambda_{k},k-1)\not\in\lambda^{\prime}/\mu^{\prime}$. Moreover, if $s_{k}=1$, we obtain (3.14) and (3.16) by Remark 3.6. Note that if $s_{k}=1$ and $\lambda_{k}=\mu_{k}+1$, then $G_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=G_{(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ because there are no cells in column $k$ of $(\lambda-\epsilon_{k})^{\prime}/\mu^{\prime}$, hence the bound condition for column $k$ is irrelevant.
For the second identity let
$E_{i,j}$ $=e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{i-1}-A_{j}-B_{\lambda_{i}}+B_{\mu_{j}})],$
$F_{i,j}$ $=e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}-\chi(i=k)]}\ominus(A_{i-1}-A_{j}-B_{\lambda_{i}}+B_{\mu_{j}})],$
$G_{i,j}$ $=e_{\lambda_{i}-\mu_{j}-i+j-\chi(i=k)}[X_{[r_{j},s_{i}-\chi(i=k)]}\ominus(A_{i-1}-A_{j}-B_{\lambda_{i}-\chi(i=k)}+B_{\mu_{j}})].$
Then it suffices to show that
(3.18) $E_{k,j}=(1-x_{s_{k}}\beta_{\lambda_{k}})F_{k,j}+x_{s_{k}}G_{k,j}$
for all $1\leq j\leq n$. In order to prove (3.18), we consider the two cases $r_{j}\leq s_{k}$ and $r_{j}>s_{k}$. Let $m=\lambda_{k}-\mu_{j}-k+j$ and $Z=A_{k-1}-A_{j}-B_{\lambda_{k}}+B_{\mu_{j}}$.
First, suppose $r_{j}\leq s_{k}$. We use the same argument in the proof of Lemma 3.15. Since $e_{m+\ell}[X_{[r_{j},s_{k}-1]}]=e_{m+\ell}[X_{[r_{j},s_{k}]}-x_{s_{k}}]$, by Lemma 2.3, we have
$e_{m+\ell}\left[X_{[r_{j},s_{k}]}\right]$ $=e_{m+\ell}\left[X_{[r_{j},s_{k}-1]}\right]+x_{s_{k}}e_{m+\ell-1}\left[X_{[r_{j},s_{k}-1]}\right],$
$e_{\ell}[Z]$ $=e_{\ell}[Z+\beta_{\lambda_{k}}]-\beta_{\lambda_{k}}e_{\ell-1}[Z].$
Thus
$E_{k,j}$ $=\sum_{\ell\geq 0}e_{m+\ell}\left[X_{[r_{j},s_{k}-1]}\right]e_{\ell}[Z]+x_{s_{k}}\sum_{\ell\geq 0}e_{m+\ell-1}\left[X_{[r_{j},s_{k}-1]}\right]e_{\ell}[Z]$
$=F_{k,j}+x_{s_{k}}\sum_{\ell\geq 0}e_{m+\ell-1}\left[X_{[r_{j},s_{k}-1]}\right]\left(e_{\ell}[Z+\beta_{\lambda_{k}}]-\beta_{\lambda_{k}}e_{\ell-1}[Z]\right)$
$=F_{k,j}+x_{s_{k}}G_{k,j}-x_{s_{k}}\beta_{\lambda_{k}}F_{k,j},$
as required.
Suppose $r_{j}>s_{k}$. Since $X_{[r_{j},s_{k}]}=X_{[r_{j},s_{k}-1]}=0$, we obtain
$E_{k,j}=F_{k,j}=e_{m}[0\ominus Z]=e_{-m}[Z],\qquad G_{k,j}=e_{1-m}[Z+\beta_{\lambda_{k}}].$
By the assumption $\mu_{k}<\lambda_{k}$ in the lemma, we have $j\neq k$ because otherwise $s_{k}<r_{j}=r_{k}\leq s_{k}+\mu_{k}-\lambda_{k}+1\leq s_{k}$, which is impossible. If $j>k$, then $\mu_{j}\leq\mu_{k}<\lambda_{k}$. Thus
$1-m=(1+\mu_{j}-\lambda_{k})+(k-j)\leq 0+(-1)<0,$
which implies $E_{k,j}=F_{k,j}=G_{k,j}=0$. Now suppose $j<k$. Since $r_{j}-\mu_{j}\leq r_{k}-\mu_{k}\leq s_{k}-\lambda_{k}+1$, we obtain $s_{k}<r_{j}\leq s_{k}+\mu_{j}-\lambda_{k}+1.$ This gives $\mu_{j}-\lambda_{k}\geq 0$ and hence
$e_{-m}[Z]=e_{-m}[A_{k-1}-A_{j}-B_{\lambda_{k}}+B_{\mu_{j}}]=e_{-m}(\alpha_{j+1},\dots,\alpha_{k-1},\beta_{\lambda_{k}+1},\dots,\beta_{\mu_{j}})=0,$
because the number of variables is $k-1-j+\mu_{j}-\lambda_{k}=-m-1<-m$ (which also implies $-m>0$). Therefore $E_{k,j}=F_{k,j}=e_{-m}[Z]=0$ and similarly we obtain $G_{k,j}=e_{1-m}[Z+\beta_{\lambda_{k}}]=0$. This shows (3.18).
To complete the proof, it remains to show (3.15) and (3.17) for the case $s_{k}=1$. Since $r_{j}>s_{k}-1=0$, (3.15) and (3.17) for the case $\lambda_{k}>\mu_{k}+1$ can be shown by applying the same argument as before. Suppose $s_{k}=1$ and $\lambda_{k}=\mu_{k}+1$. Then, since $\mu_{k}=(\lambda-\epsilon_{k})_{k}$, we have $\mu_{k-1}\geq\mu_{k}=(\lambda-\epsilon_{k})_{k}$ and $\mu_{k}=(\lambda-\epsilon_{k})_{k}\geq(\lambda-\epsilon_{k})_{k+1}$. Therefore, by applying Lemma 3.18 twice, we obtain
$\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{({\lambda}^{(1)})^{\prime}/({\mu}^{(1)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\\ \times\phi_{\bm{\alpha}}^{k-1}\left(\tilde{G}_{((\lambda_{k}-1))^{\prime}/((\mu_{k}))^{\prime}}^{\operatorname{col}((r_{k}),(s_{k}-1))}(\bm{x};\bm{\alpha},\bm{\beta})\right)\phi_{\bm{\alpha}}^{k}\left(\tilde{G}_{({\lambda}^{(2)})^{\prime}/({\mu}^{(2)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right),$
where $\bm{r}^{(1)}=(r_{1},\dots,r_{k-1})$ and $\bm{r}^{(2)}=(r_{k+1},\dots,r_{n})$, and $\bm{s}^{(1)}$, $\bm{s}^{(2)}$, $\lambda^{(1)}$, $\lambda^{(2)}$, $\mu^{(1)}$, and $\mu^{(2)}$ are defined similarly. Since $(\lambda_{k}-1)=(\mu_{k})$ as partitions, by Lemma 3.19, the above identity can be rewritten as
(3.19) $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{({\lambda}^{(1)})^{\prime}/({\mu}^{(1)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(1)},{\bm{s}}^{(1)})}(\bm{x};\bm{\alpha},\bm{\beta})\phi_{\bm{\alpha}}^{k}\left(\tilde{G}_{({\lambda}^{(2)})^{\prime}/({\mu}^{(2)})^{\prime}}^{\operatorname{col}({\bm{r}}^{(2)},{\bm{s}}^{(2)})}(\bm{x};\bm{\alpha},\bm{\beta})\right).$
The same argument also shows that $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ is equal to the right-hand side of (3.19). Therefore we obtain (3.17). ∎
Lemma 3.22.
Let $\lambda,\mu\in\operatorname{Par}_{n}$ and $\bm{r},\bm{s}\in\mathbb{P}^{n}$. Suppose $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}$ for all $11\leq i\leq n-$ 1. Let $11\leq k\leq n-$ 1 be an integer satisfying the following conditions:
  1. (1)
    $\mu_{k}<\lambda_{k}$,
  2. (2)
    $r_{k}-\mu_{k}\leq s_{k}-\lambda_{k}+1$,
  3. (3)
    $s_{k}-1=s_{k+1}$, and
  4. (4)
    $\lambda_{k}=\lambda_{k+1}$.
Then
$G_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=G_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta}),$
$\tilde{G}_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\tilde{G}_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Moreover, if $s_{k}=1$, then
$G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0.$
Proof.
The proof is similar to that of Lemma 3.21. Let $T\in\operatorname{MMSVT}^{\operatorname{col}(\bm{r},\bm{s})}(\lambda^{\prime}/\mu^{\prime})$. Then
$\max(T(\lambda_{k},k))\leq\min(T(\lambda_{k},k+1))=\min(T(\lambda_{k+1},k+1))\leq s_{k+1}=s_{k}-1.$
Thus every integer in the $k$th column of $T$ is at most $s_{k}-1$, which shows the first identity. As before, we have $G_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0$ if $s_{k}=1$ by Remark 3.6.
For the second identity let
$E_{i,j}$ $=e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}\ominus(A_{i-1}-A_{j}-B_{\lambda_{i}}+B_{\mu_{j}})],$
$F_{i,j}$ $=e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}-\chi(i=k)]}\ominus(A_{i-1}-A_{j}-B_{\lambda_{i}}+B_{\mu_{j}})].$
Then it suffices to show that
(3.20) $E_{k,j}=(1-x_{s_{k}}\alpha_{k})F_{k,j}+x_{s_{k}}F_{k+1,j}$
for all $1\leq j\leq n$. In order to prove (3.20), we consider the two cases $r_{j}\leq s_{k}$ and $r_{j}>s_{k}$. Let $m=\lambda_{k}-\mu_{j}-k+j$ and $Z=A_{k-1}-A_{j}-B_{\lambda_{k}}+B_{\mu_{j}}$.
First, suppose $r_{j}\leq s_{k}$. By Lemma 2.3, we have
$e_{m+\ell}\left[X_{[r_{j},s_{k}]}\right]$ $=e_{m+\ell}\left[X_{[r_{j},s_{k}-1]}\right]+x_{s_{k}}e_{m+\ell-1}\left[X_{[r_{j},s_{k}-1]}\right],$
$e_{\ell}[Z]$ $=e_{\ell}[Z+\alpha_{k}]-\alpha_{k}e_{\ell-1}[Z].$
Thus
$E_{k,j}$ $=\sum_{\ell\geq 0}e_{m+\ell}\left[X_{[r_{j},s_{k}-1]}\right]e_{\ell}[Z]+x_{s_{k}}\sum_{\ell\geq 0}e_{m+\ell-1}\left[X_{[r_{j},s_{k}-1]}\right]e_{\ell}[Z]$
$=F_{k,j}+x_{s_{k}}\sum_{\ell\geq 0}e_{m+\ell-1}\left[X_{[r_{j},s_{k}-1]}\right]\left(e_{\ell}[Z+\alpha_{k}]-\alpha_{k}e_{\ell-1}[Z]\right)$
$=F_{k,j}+x_{s_{k}}F_{k+1,j}-x_{s_{k}}\alpha_{k}F_{k,j},$
as required.
Suppose $r_{j}>s_{k}$. Since $X_{[r_{j},s_{k}]}=X_{[r_{j},s_{k}-1]}=0$, we obtain
$E_{k,j}=F_{k,j}=e_{m}[0\ominus Z]=e_{-m}[Z],\qquad F_{k+1,j}=e_{m-1}[0\ominus Z]=e_{1-m}[Z].$
Then by the same argument as in the proof of Lemma 3.21, we obtain $E_{k,j}=F_{k,j}=F_{k+1,j}=0$, which shows (3.20).
Finally, if $s_{k}=1$, then $r_{j}>s_{k}-1=0$ and hence we obtain $F_{k,j}=0$ by the argument above. This shows $\tilde{G}_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s}-\epsilon_{k})}(\bm{x};\bm{\alpha},\bm{\beta})=0$ for $s_{k}=1$, which completes the proof. ∎
We are ready to prove Theorem 3.8, which is restated as follows.
Theorem 3.23.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. If $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}$ and $s_{i}-\lambda_{i}\leq s_{i+1}-\lambda_{i+1}+1$ whenever $\mu_{i}<\lambda_{i+1}$ for $11\leq i\leq n-$ 1, then
$G_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=\tilde{G}_{{\lambda^{\prime}/\mu^{\prime}}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta}).$
Proof.
The proof is similar to that of Theorem 3.17. We proceed by strong induction on $N:=n+|{\lambda/\mu}|+\sum_{i=1}^{n}s_{i}$. Suppose that the theorem holds for all integers less than $N$. If $\mu=\lambda$, then it follows from Lemma 3.19. Therefore we may assume $\mu\neq\lambda$. We consider the following two cases.
Case 1: Suppose that there is an integer $11\leq t\leq n-$ 1 satisfying $\mu_{t}\geq\lambda_{t+1}$. We set $\gamma^{(1)}=(\gamma_{1},\dots,\gamma_{t})$ and $\gamma^{(2)}=(\gamma_{t+1},\dots,\gamma_{n})$ for each $\gamma\in\{\lambda,\mu,\bm{r},\bm{s}\}$. By Lemma 3.18, the statement is separated into the cases $(\lambda^{(i)},\mu^{(i)},\bm{r}^{(i)},\bm{s}^{(i)})$ for $i=1,2$. Each case is then covered by the induction hypothesis.
Case 2: Suppose that $\mu_{t}<\lambda_{t+1}$ for all $11\leq t\leq n-$ 1 (or $n=1$). Then by the assumption, we have $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}$ and $s_{i}-\lambda_{i}\leq s_{i+1}-\lambda_{i+1}+1$ for all $11\leq i\leq n-$ 1. Note also that $\mu_{i}<\lambda_{i}$ for all $1\leq i\leq n$ since $\mu_{i}<\lambda_{i+1}\leq\lambda_{i}$ for $11\leq i\leq n-$ 1 and $\mu_{n}\leq\mu_{n-1}<\lambda_{n}$.
Now let
$k=\max\left(\{1\}\cup\{2\leq m\leq n:s_{m-1}-\lambda_{m-1}\leq s_{m}-\lambda_{m}\}\right).$
Observe that, by the definition of $k$, we have $s_{k-1}-\lambda_{k-1}\leq s_{k}-\lambda_{k}$ if $k\geq 2$. Moreover, since $s_{i}-\lambda_{i}\leq s_{i+1}-\lambda_{i+1}+1$ for all $11\leq i\leq n-$ 1, we have
(3.21) $s_{k}-\lambda_{k}=s_{k+1}-\lambda_{k+1}+1\qquad\mbox{if $k<n$.}$
If $s_{k}-\lambda_{k}+1<r_{k}-\mu_{k}$, then the statement holds by Lemma 3.20. Hence we may assume $r_{k}-\mu_{k}\leq s_{k}-\lambda_{k}+1$. If $\lambda_{k}>\lambda_{k+1}$, then we apply Lemma 3.21. If $\lambda_{k}=\lambda_{k+1}$, then $k<n$ because otherwise $00\leq\mu_{n}<\lambda_{n}=\lambda_{n+1}=$ 0, which is a contradiction. In this case, by (3.21), we have $s_{k}=s_{k+1}-\lambda_{k+1}+\lambda_{k}+1=s_{k+1}+1$. Thus we can apply Lemma 3.22. In either case the statement reduces to some cases with a smaller $N$. Thus, by the induction hypothesis, we obtain the statement.
The above two cases show that the statement holds for $N$, and the theorem follows by induction. ∎
Remark 3.24.
Computer experiments suggest that the condition $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}$ in Theorem 3.8 may be replaced by the weaker condition $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}+1$ as in a result of Wachs [19, Theorem 3.5*]. Under this weaker condition, the inequality $\mu_{j}-\lambda_{k}\geq 0$ does not hold when $j<k$ in the last part of the proof of Lemma 3.21. Thus we must also consider the case $\mu_{j}<\lambda_{k}$ when $j<k$ to complete the proof of Lemma 3.21 under the weaker condition. All other lemmas in this subsection containing the condition $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}$ can be proved similarly under the weaker condition $r_{i}-\mu_{i}\leq r_{i+1}-\mu_{i+1}+1$.
3.3. Matsumura’s flagged Grothendieck polynomials
Matsumura [17] found a Jacobi–Trudi-like formula for flagged Grothendieck polynomials. In this subsection we show that our Jacobi–Trudi-like formula implies his result.
Following the notation in [17], for $\lambda,\mu\in\operatorname{Par}_{n}$ and $f=(f_{1},\dots,f_{n}),g=(g_{1},\dots,g_{n})\in\mathbb{P}^{n}$, let
$G_{\lambda/\mu,f/g}(\bm{x}):=\sum_{T\in\operatorname{FSVT}(\lambda/\mu,f/g)}\beta^{|T|-|{\lambda/\mu}|}\bm{x}^{T},$
where $\operatorname{FSVT}(\lambda/\mu,f/g)$ is the set of set-valued tableaux $T$ of shape ${\lambda/\mu}$ such that every integer appearing in row $i$ is in $\{g_{i},g_{i}+1,\dots,f_{i}\}$, and $|T|$ is the total number of integers in $T$.
We define $G^{[p/q]}_{m}(\bm{x})$ by the following Laurent series in $u$:
(3.22) $\sum_{m\in\mathbb{Z}}G^{[p/q]}_{m}(\bm{x})u^{m}=\frac{1}{1+\beta u^{-1}}\prod_{q\leq i\leq p}\frac{1+\beta x_{i}}{1-x_{i}u},$
where
$\frac{1}{1+\beta u^{-1}}=\frac{1}{1-(-\beta u^{-1})}=1+(-\beta u^{-1})+(-\beta u^{-1})^{2}+\cdots,$
and
$\frac{1+\beta x_{i}}{1-x_{i}u}=(1+\beta x_{i})(1+x_{i}u+x_{i}^{2}u^{2}+\cdots).$
If $p<q$, then the product in (3.22) is defined to be $1$. We note that the definition of $G^{[p/q]}_{m}(\bm{x})$ in [17, p.224] has a typo, $\beta^{-1}u$, which should be $\beta u^{-1}$.
We now state Matsumura’s theorem.
Theorem 3.25.
[17, Theorem 4] Let $\lambda,\mu\in\operatorname{Par}_{n}$ and $f=(f_{1},\dots,f_{n}),g=(g_{1},\dots,g_{n})\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$ and $g_{i}\leq g_{i+1},f_{i}\leq f_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$. We have
(3.23) $G_{\lambda/\mu,f/g}(\bm{x})=\det\left(\sum_{s\geq 0}\binom{i-j}{s}\beta^{s}G^{[f_{i}/g_{j}]}_{\lambda_{i}-\mu_{j}-i+j+s}(\bm{x})\right)_{i,j=1}^{\ell(\lambda)}.$
In what follows we reformulate Theorem 3.25 using our notation $\ominus$. Recall that $h_{n}(\bm{x})=0$ for $n<0$.
Lemma 3.26.
For any integer $m$ and any positive integers $p$ and $q$, we have
$G^{[p/q]}_{m}(\bm{x})=\prod_{i=q}^{p}(1+\beta x_{i})\sum_{k\geq 0}(-\beta)^{k}h_{m+k}[X_{[q,p]}].$
Proof.
Since $h_{n}(\bm{x})=0$ for $n<0$, we have
$\prod_{q\leq i\leq p}\frac{1}{1-x_{i}u}=\sum_{l\geq 0}h_{l}[X_{[q,p]}]u^{l}=\sum_{l\in\mathbb{Z}}h_{l}[X_{[q,p]}]u^{l}.$
Thus the right hand side of (3.22) is
$\frac{1}{1+\beta u^{-1}}\prod_{q\leq i\leq p}\frac{1+\beta x_{i}}{1-x_{i}u}$ $=\prod_{i=q}^{p}(1+\beta x_{i})\sum_{k\geq 0}(-\beta)^{k}u^{-k}\sum_{l\in\mathbb{Z}}h_{l}[X_{[q,p]}]u^{l}$
$=\prod_{i=q}^{p}(1+\beta x_{i})\sum_{m\in\mathbb{Z}}\left(\sum_{k\geq 0}(-\beta)^{k}h_{m+k}[X_{[q,p]}]\right)u^{m},$
which shows the lemma. ∎
Using Lemma 3.26, the $(i,j)$-entry of the matrix in (3.23) can be rewritten as follows. Here, to avoid confusion, we write $\overline{\beta}=-\beta$ as a plethystic variable.
Lemma 3.27.
We have
$\sum_{s\geq 0}\binom{i-j}{s}\beta^{s}G^{[f_{i}/g_{j}]}_{\lambda_{i}-\mu_{j}-i+j+s}(\bm{x})=C_{i,j}h_{\lambda_{i}-\mu_{j}-i+j}[X_{[g_{j},f_{i}]}\ominus(j-i+1)\overline{\beta}],$
where $C_{i,j}=\prod_{l=g_{j}}^{f_{i}}(1+\beta x_{l})$.
Proof.
By Lemma 3.26, we have
$\sum_{s\geq 0}\binom{i-j}{s}\beta^{s}G^{[f_{i}/g_{j}]}_{\lambda_{i}-\mu_{j}-i+j+s}(\bm{x})$ $=C_{i,j}\sum_{s\geq 0}\binom{i-j}{s}\beta^{s}\sum_{k\geq 0}(-\beta)^{k}h_{\lambda_{i}-\mu_{j}-i+j+s+k}[X_{[g_{j},f_{i}]}]$
$=C_{i,j}\sum_{m\geq 0}\beta^{m}h_{\lambda_{i}-\mu_{j}-i+j+m}[X_{[g_{j},f_{i}]}]\sum_{s=0}^{m}\binom{i-j}{s}(-1)^{m-s}$
(3.24) $=C_{i,j}\sum_{m\geq 0}\beta^{m}h_{\lambda_{i}-\mu_{j}-i+j+m}[X_{[g_{j},f_{i}]}]\binom{i-j-1}{m},$
where the identity $\sum_{s=0}^{m}\binom{n}{s}(-1)^{m-s}=\binom{n-1}{m}$ is used. We claim that
(3.25) $\beta^{m}\binom{i-j-1}{m}=h_{m}[(j-i+1)\overline{\beta}].$
Note that, by (3.25), the right side of (3.24) is equal to
$C_{i,j}\sum_{m\geq 0}h_{\lambda_{i}-\mu_{j}-i+j+m}[X_{[g_{j},f_{i}]}]h_{m}[(j-i+1)\overline{\beta}]=C_{i,j}h_{\lambda_{i}-\mu_{j}-i+j}[X_{[g_{j},f_{i}]}\ominus(j-i+1)\overline{\beta}],$
which implies the lemma. Thus it remains to prove (3.25).
If $i>j$, then
$\beta^{m}\binom{i-j-1}{m}=(-1)^{m}(-\beta)^{m}\binom{i-j-1}{m}=(-1)^{m}e_{m}[(i-j-1)\overline{\beta}]=h_{m}[(j-i+1)\overline{\beta}].$
Now suppose $i\leq j$. Then, by the identity $(-1)^{m}\binom{-n}{m}=\left(\!\binom{n}{m}\!\right):=\binom{n+m-1}{m}$, we have
$\beta^{m}\binom{i-j-1}{m}=(-\beta)^{m}(-1)^{m}\binom{i-j-1}{m}=(-\beta)^{m}\left(\!\binom{j-i+1}{m}\!\right)=h_{m}[(j-i+1)\overline{\beta}].$
Hence we always have (3.25) and the proof is completed. ∎
Using Lemma 3.27 we restate Theorem 3.25 as follows.
Theorem 3.28 (Matsumura’s theorem reformulated).
Let $\lambda,\mu\in\operatorname{Par}_{n}$ and $f=(f_{1},\dots,f_{n}),g=(g_{1},\dots,g_{n})\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$ and $g_{i}\leq g_{i+1},f_{i}\leq f_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$. We have
(3.26) $G_{\lambda/\mu,f/g}(\bm{x})=C_{\ell(\lambda)}\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[X_{[g_{j},f_{i}]}\ominus(j-i+1)\overline{\beta}]\right)_{i,j=1}^{\ell(\lambda)},$
where $\overline{\beta}=-\beta$ and $C_{\ell(\lambda)}=\prod_{i=1}^{\ell(\lambda)}\prod_{l=g_{i}}^{f_{i}}(1+\beta x_{l})$.
[…]
4. Flagged dual Grothendieck polynomials
The structure of this section is similar to that of the previous section. In this section we give a combinatorial model for the refined dual canonical stable Grothendieck polynomials $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ using marked reverse plane partitions. To this end we introduce a flagged version of $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ using marked reverse plane partitions and prove a Jacobi–Trudi-like formula for this flagged version, which reduces to the Jacobi–Trudi-like formula for $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ in Theorem 1.2. More generally, we consider two flagged versions of $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$ and extend the partition $\lambda$ to a skew shape.
Definition 4.1.
A reverse plane partition of shape $\lambda/\mu$ is a filling $T$ of the cells in $\lambda/\mu$ with positive integers such that the integers are weakly increasing in each row and in each column. A left-marked reverse plane partition (or simply marked reverse plane partition) of shape ${\lambda/\mu}$ is a reverse plane partition $T$ of shape ${\lambda/\mu}$ in which every entry $T(i,j)$ with $T(i,j)=T(i,j+1)$ may be marked. We denote by ${\operatorname{MRPP}}({\lambda/\mu})$ the set of marked reverse plane partitions of shape ${\lambda/\mu}$. For $T\in{\operatorname{MRPP}}({\lambda/\mu})$, define
$\operatorname{wt}(T)=\prod_{(i,j)\in{\lambda/\mu}}\operatorname{wt}(T(i,j)),$
where
$\operatorname{wt}(T(i,j))=\begin{cases}-\alpha_{j}&\mbox{if $T(i,j)$ is marked,}\\ \beta_{i-1}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i-1,j)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i-1,j)$.}\end{cases}$
Here, the equality $T(i,j)=T(i-1,j)$ means that their underlying integers are equal. See Figure 3.
[…]
Figure 3. An example of $T\in{\operatorname{MRPP}}({\lambda/\mu})$, where $\lambda=(6,5,3,3)$ and $\mu=(2,1,1)$. We have the weight $\operatorname{wt}(T)=x_{1}x_{2}x_{3}^{2}x_{4}x_{5}(-\alpha_{1})(-\alpha_{2})^{2}(-\alpha_{5})\beta_{1}\beta_{2}^{2}$.
The main goal of this section is to prove the following combinatorial interpretation for $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$.
Theorem 4.2.
We have
(4.1) $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})=\sum_{T\in{\operatorname{MRPP}}(\lambda)}\operatorname{wt}(T).$
Remark 4.3.
[…]
Similar to our approach in the previous section, in order to prove Theorem 4.2 we introduce a flagged version. For completeness we again consider row- and column-flagged versions on skew shapes.
Definition 4.4.
Let $\bm{r}=(r_{1},\dots,r_{n})\in\mathbb{P}^{n}$ and $\bm{s}=(s_{1},\dots,s_{n})\in\mathbb{P}^{n}$. We denote by ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})$ the set of $T\in{\operatorname{MRPP}}({\lambda/\mu})$ such that $r_{i}\leq T(i,j)\leq s_{i}$ for all $(i,j)\in{\lambda/\mu}$. Similarly, ${\operatorname{MRPP}}^{\operatorname{col}(\bm{r},\bm{s})}({\lambda/\mu})$ denotes the set of $T\in{\operatorname{MRPP}}({\lambda/\mu})$ such that $r_{j}\leq T(i,j)\leq s_{j}$ for all $(i,j)\in{\lambda/\mu}$. We define row-flagged and column-flagged refined dual canonical stable Grothendieck polynomials by
$g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\sum_{T\in{\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})}\operatorname{wt}(T),$
$g_{\lambda/\mu}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\sum_{T\in{\operatorname{MRPP}}^{\operatorname{col}(\bm{r},\bm{s})}({\lambda/\mu})}\operatorname{wt}(T).$
Now we state Jacobi–Trudi-like formulas for the two flagged versions of $g_{\lambda}(\bm{x};\bm{\alpha},\bm{\beta})$.
Theorem 4.5.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ such that $r_{i}\leq r_{i+1}$ and $s_{i}\leq s_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$. Then we have
(4.2) $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{\lambda_{i}-1}+A_{\mu_{j}}+B_{i-1}-B_{j-1}]\right)_{i,j=1}^{n},$
(4.3) $g_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})$ $=\det\left(e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1}-B_{\mu_{j}}]\right)_{i,j=1}^{n}.$
The proof of this theorem will be given in the next two subsections. Our proof is based on the ideas in [11]. We introduce new notions in Definitions 4.13 and 4.23, which simplify the proofs in [11].
By specializing Theorem 4.5 we obtain the following corollary, which implies the combinatorial interpretation for $g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})$ in Theorem 4.2 as $n\to\infty$.
Corollary 4.6.
For a partition $\lambda\in\operatorname{Par}_{n}$, we have
$g_{\lambda}(\bm{x}_{n};\bm{\alpha},\bm{\beta})=\sum_{T\in{\operatorname{MRPP}}(\lambda),\max(T)\leq n}\operatorname{wt}(T),$
where $\max(T)$ is the largest entry in $T$.
Proof.
This can be proved by the same argument as in the proof of Corollary 3.10 using (4.2) and Theorem 1.2. ∎
Remark 4.7.
The identity (4.3) with $\bm{\alpha}=(0,0,\dots)$ is proved in [11, Theorem 1.4] under the weaker assumption $r_{i}\leq r_{i+1}+1$ and $s_{i}\leq s_{i+1}+1$ whenever $\mu_{i}<\lambda_{i+1}$. We cannot replace the assumption in Theorem 4.5 by this weaker one. For example, if $\lambda=(1,1),\mu=(0,0),\bm{r}=(1,1)$ and $\bm{s}=(2,1)$, then the left and right sides of (4.3) are $x_{1}^{2}-\alpha_{1}x_{1}$ and $x_{1}^{2}-\alpha_{1}x_{1}-\alpha_{1}x_{2}$, respectively.
The following lemmas show that we may assume $\mu\subseteq\lambda$ and $\bm{r}\leq\bm{s}$ when we prove Theorem 4.5.
Lemma 4.8.
Under the assumption in Theorem 4.5, if $\mu\not\subseteq\lambda$, then the both sides of (4.2) and (4.3) are equal to 00.
Proof.
By definition, if $\mu\not\subseteq\lambda$, we have $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=g_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=0$. By Lemma 2.4, the determinants in (4.2) and (4.3) are equal to 00. ∎
Lemma 4.9.
Under the assumption in Theorem 4.5, if $\mu_{k}<\lambda_{k}$ and $r_{k}>s_{k}$ for some $k$, then the both sides of (4.2) and (4.3) are equal to 00.
Proof.
Since $\mu_{k}<\lambda_{k}$ and $r_{k}>s_{k}$, we have ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})={\operatorname{MRPP}}^{\operatorname{col}(\bm{r},\bm{s})}(\lambda^{\prime}/\mu^{\prime})=\emptyset$ and therefore $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=g_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})=0$.
We now show that the right hand side of (4.2) and (4.3) are equal to 00. By Lemma 4.8 and Lemma 2.5, we may assume $\mu_{m}<\lambda_{m}$ for all $m$. Let $(R_{i,j})_{i,j=1}^{n}$ be the matrix in the right hand side of (4.2).
We first suppose that $\mu_{m}<\lambda_{m+1}$ for all $m$. Then the assumption in Theorem 4.5 gives $r_{m}\leq r_{m+1}$ and $s_{m}\leq s_{m+1}$ for all $m$. Thus, for $1\leq i\leq k$ and $k\leq j\leq n$, we have $r_{j}\geq r_{k}>s_{k}\geq s_{i}$ and $\mu_{j}\leq\mu_{k}<\lambda_{k}\leq\lambda_{i}$, which imply
$R_{i,j}$ $=h_{\lambda_{i}-\mu_{j}-i+j}[-A_{\lambda_{i}-1}+A_{\mu_{j}}+B_{i-1}-B_{j-1}]$
$=h_{\lambda_{i}-\mu_{j}-i+j}[-(\alpha_{\mu_{j}+1}+\dots+\alpha_{\lambda_{i}-1})-(\beta_{i}+\dots+\beta_{j-1})]$
(4.4) $=(-1)^{\lambda_{i}-\mu_{j}-i+j}e_{\lambda_{i}-\mu_{j}-i+j}(\alpha_{\mu_{j}+1},\dots,\alpha_{\lambda_{i}-1},\beta_{i},\dots,\beta_{j-1}).$
Since the number of variables in (4.4) is less than $\lambda_{i}-\mu_{j}-i+j>0$, we have $R_{i,j}=0$ for all $1\leq i\leq k$ and $k\leq j\leq n$. This implies that the determinant in (4.2) is 00.
Now we suppose $\mu_{m}\geq\lambda_{m+1}$ for some $m$. For $m+1\leq i\leq n$ and $1\leq j\leq m$, we have $\mu_{j}\geq\mu_{m}\geq\lambda_{m+1}\geq\lambda_{i}$ and therefore $\lambda_{i}-\mu_{j}-i+j<0$. Then, since $R_{i,j}=h_{\lambda_{i}-\mu_{j}-i+j}[Z]$ for some formal power series $Z$, we have $R_{i,j}=0$. Thus the matrix $(R_{i,j})_{i,j=1}^{n}$ is a block upper triangular matrix and its determinant is written as
$\det(R_{i,j})_{i,j=1}^{n}=\det(R_{i,j})_{i,j=1}^{m}\det(R_{i,j})_{i,j=m+1}^{n}.$
Therefore, one can inductively deduce that the determinant in (4.2) is equal to 00.
Similarly, one can show that the right hand side of (4.3) is equal to 00. ∎
The rest of this section is devoted to proving Theorem 4.5.
4.1. Proof of (4.2): row-flagged dual Grothendieck polynomials
In this subsection we prove the first identity (4.2) in Theorem 4.5. Our strategy is as follows. We first show that (4.2) is equivalent to the following formula:
(4.5) $g_{\lambda/\mu}^{\operatorname{row}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})\\ =\det\left(\chi(r_{j}\leq s_{i})h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{\lambda_{i}-1}+A_{\mu_{j}}+B_{i-1}-B_{j-1}]\right)_{i,j=1}^{n}.$
We then find a generalization of this new formula and prove it using recursions.
The following lemma shows that (4.2) and (4.5) are equivalent.
Lemma 4.10.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$ and $\bm{r}\leq\bm{s}$ such that $r_{i}\leq r_{i+1}$ and $s_{i}\leq s_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$. Then
$\det\left(\chi(r_{j}\leq s_{i})h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{\lambda_{i}-1}+A_{\mu_{j}}+B_{i-1}-B_{j-1}]\right)_{i,j=1}^{n}\\ =\det\left(h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{\lambda_{i}-1}+A_{\mu_{j}}+B_{i-1}-B_{j-1}]\right)_{i,j=1}^{n}.$
Proof.
Let $L=(L_{i,j})_{i,j=1}^{n}$ and $R=(R_{i,j})_{i,j=1}^{n}$ be the matrices in the left and right sides of the equation, respectively. If $r_{j}\leq s_{i}$ for all $i$ and $j$, then the matrices $L$ and $R$ are the same, so it is done. Assume that $L_{i,j}\neq R_{i,j}$ for some $i$ and $j$. In this case we must have $r_{j}>s_{i}$ and $L_{i,j}=0$. Thus it is enough to show that the $(i,j)$-entry $R_{i,j}$ does not contribute to the determinant of the matrix $R$. We consider the two cases $i>j$ and $i<j$.
First suppose $i>j$. Since $r_{i}\leq s_{i}<r_{j}$, by the condition that $r_{k}\leq r_{k+1}$ whenever $\mu_{k}<\lambda_{k+1}$, there is at least one $m\in\{j,\dots,i-1\}$ such that $\mu_{m}\geq\lambda_{m+1}$. Then for integers $p$ and $q$ satisfying $1\leq p\leq m<q\leq n$, we have $\lambda_{q}-\mu_{p}+p-q\leq\lambda_{m+1}-\mu_{m}+m-(m+1)<0$, which implies $R_{p,q}=0$. Therefore, $\det R=\det(R_{k,l})_{k,l=1}^{m}\det(R_{k,l})_{k,l=m+1}^{n}$. Since $j\leq m<i$, this means that the $(i,j)$-entry $R_{i,j}$ does not contribute to the determinant of the matrix $R$.
We now suppose $i<j$. If $\mu_{j}<\lambda_{i}$, then by the same argument in (4.4), we have $R_{i,j}=0$. This is a contradiction to the assumption $L_{i,j}\neq R_{i,j}$, so we must have $\lambda_{i}\leq\mu_{j}$. Then, since $\mu_{i}\leq\lambda_{i}\leq\mu_{j}\leq\mu_{i}$, we also have $\mu_{i}=\lambda_{i}$. For integers $p$ and $q$ satisfying $1\leq p\leq i\leq q\leq n$, we have $\lambda_{q}-\mu_{p}+p-q\leq 0$ and the equality holds if and only if $p=q=i$, which means that $L_{q,p}=R_{q,p}=\chi(p=q=i)$. From this, the matrix $R$ is a block upper triangular matrix and it gives $\det R=\det(R_{k,l})_{k,l=1}^{i-1}\det(R_{k,l})_{k,l=i+1}^{n}$. This means that the $(i,j)$-entry of $R$ does not contribute to the determinant of the matrix $R$, which completes the proof. ∎
Our next step is to find a generalization of (4.5). To do this we need the following definitions.
Definition 4.11.
A dented partition is a sequence $(\lambda_{1},\dots,\lambda_{n})$ of nonnegative integers such that
$\lambda_{1}+1=\dots=\lambda_{k-1}+1=\lambda_{k}\geq\dots\geq\lambda_{n},$
for some $1\leq k\leq n$. The cell $(k,\lambda_{k})$ is called the minimal cell of $\lambda$. Similar to partitions, each $\lambda_{i}$ of a dented partition $\lambda=(\lambda_{1},\dots,\lambda_{n})$ is called a part. The Young diagram of $\lambda$ is also defined as usual. We denote by $\operatorname{DPar}_{n}$ the set of all dented partitions with at most $n$ parts. See Figure 4.
[…]
boxsize=1.5em {ytableau}   &
      *(gray)
Figure 4. The Young diagram of a dented partition $\lambda=(3,3,4,4,1)$. The cell $(3,4)$ is the minimal cell, which is colored gray.
Note that a dented partition $\lambda$ is a partition if and only if its minimal cell is $(1,\lambda_{1})$.
Remark 4.12.
The motivation of introducing dented partitions is to construct a marked RPP of shape $\lambda$ by filling a cell one at a time with respect to the total order $\prec$ on the cells defined by $(i,j)\prec(i^{\prime},j^{\prime})$ if $j>j^{\prime}$ or $j=j^{\prime}$ and $i\leq i^{\prime}$. At each stage of the construction, the remaining cells form a dented partition and the next cell to be filled is the minimal cell of this dented partition.
We also note that dented partitions are a special case of pseudo-partitions introduced by Aas, Grinberg, and Scrimshaw, who found a Jacobi–Trudi-like formula for the generating function for semistandard Young tableaux of a pseudo-partition shape with a flag condition in [1, Theorem 5.3].
Definition 4.13.
Let $\lambda$ be a dented partition and $\mu$ a partition with $\mu\subseteq\lambda$. Let $\bm{r},\bm{s}\in\mathbb{P}^{n}$ and $I\subseteq[n]$. If $\bm{r}\not\leq\bm{s}$, we define ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}(\lambda/\mu)=\emptyset$. If $\bm{r}\leq\bm{s}$, we define ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$ to be the set of RPPs $T$ of shape ${\lambda/\mu}$ in which every entry $T(i,j)$ with $T(i,j)=T(i,j+1)$ may be marked, where for each $i\in[n]$ we define
$T(i,\lambda_{i}+1)=\begin{cases}\infty&\mbox{if $i\not\in I$},\\ s_{i}&\mbox{if $i\in I$.}\end{cases}$
For $T\in{\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$, we define
$\operatorname{wt}(T)=\prod_{(i,j)\in{\lambda/\mu}}\operatorname{wt}(T(i,j)),$
where
$\operatorname{wt}(T(i,j))=\begin{cases}-\alpha_{j}&\mbox{if $T(i,j)$ is marked,}\\ \beta_{i-1}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i-1,j)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i-1,j)$.}\end{cases}$
Note that in the above definition the assumption $T(i,\lambda_{i}+1)=s_{i}$ for $i\in I$ also plays a role in the definition of $\operatorname{wt}(T)$. See Figure 5. Note also that ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})=\emptyset$ if $r_{i}>s_{i}$ for some $i$ (even for the case $\mu_{i}=\lambda_{i}$). Thus we have
${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}(\mu/\mu)=\begin{cases}\emptyset&\mbox{if $\bm{r}\not\leq\bm{s}$},\\ \{\emptyset\}&\mbox{otherwise.}\end{cases}$
[…]
Figure 5. An example of $T\in{\operatorname{MRPP}}_{I}^{\operatorname{row}(\bm{r},\bm{s})}({\lambda/\mu})$, where $\lambda=(3,4,4,1),\mu=(1,1),I=\{1,3\},\bm{r}=(1,1,2,2),$ and $\bm{s}=(3,3,4,5)$. Note that $T(2,4)=3$ cannot be marked because $2\notin I$ and its contribution to the weight is $\beta_{1}$ since $1\in I$ and $T(2,4)=T(1,4)=3$. Thus we have $\operatorname{wt}(T)=x_{1}x_{2}^{3}x_{4}(-\alpha_{3})(-\alpha_{4})\beta_{1}^{2}\beta_{2}$.
The following proposition is the promised generalization of (4.5).
Proposition 4.14.
Let $\lambda\in\operatorname{DPar}_{n},\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$ such that $r_{i}\leq r_{i+1}$ and $s_{i}\leq s_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$. Suppose that $(k,\lambda_{k})$ is the minimal cell of $\lambda$ and $I=\{1,2,\dots,p\}$ or $I=\{1,2,\dots,p\}\setminus\{k\}$ for some integer $k\leq p\leq\ell(\lambda)$ satisfying
$\lambda_{1}+1=\dots=\lambda_{k-1}+1=\lambda_{k}=\lambda_{k+1}=\dots=\lambda_{p}.$
Then
(4.6) $\sum_{T\in{\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})}\operatorname{wt}(T)=\det\left(h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)\right)_{i,j=1}^{n},$
where
$h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)=\chi(r_{j}\leq s_{i})h_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{\lambda_{i}-1+\chi(i\in I)}+A_{\mu_{j}}+B_{i-1}-B_{j-1}].$
Observe that (4.5) is the special case, $\lambda\in\operatorname{Par}_{n}$ and $I=\emptyset$, of Proposition 4.14. For the rest of this subsection we prove this proposition by finding the same recursions and initial conditions for both sides of (4.6). Let
$M(\lambda,\mu,\bm{r},\bm{s},I)$ $=\sum_{T\in{\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})}\operatorname{wt}(T),$
$D(\lambda,\mu,\bm{r},\bm{s},I)$ $=\det\left(h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)\right)_{i,j=1}^{n}.$
The following two lemmas give the same initial conditions for $M(\lambda,\mu,\bm{r},\bm{s},I)$ and $D(\lambda,\mu,\bm{r},\bm{s},I)$.
Lemma 4.15.
Let $\lambda\in\operatorname{DPar}_{n},\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\operatorname{RPar}_{n}$ with $\mu\subseteq\lambda$. If $\bm{r}\not\leq\bm{s}$, then
$M(\lambda,\mu,\bm{r},\bm{s},I)=D(\lambda,\mu,\bm{r},\bm{s},I)=0.$
Proof.
By assumption there is an integer $1\leq k\leq n$ with $r_{k}>s_{k}$. By definition we have $M(\lambda,\mu,\bm{r},\bm{s},I)=0$. To show $D(\lambda,\mu,\bm{r},\bm{s},I)=0$ it suffices to show that $h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)=0$ for all $1\leq i\leq k$ and $k\leq j\leq n$. Since $\bm{r},\bm{s}\in\operatorname{RPar}_{n}$, we have $r_{j}\geq r_{k}>s_{k}\geq s_{i}$, and therefore $\chi(r_{j}\leq s_{i})=0$. This implies $h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)=0$. ∎
Lemma 4.16.
Let $\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$. If $\bm{r}\leq\bm{s}$, then
$M(\mu,\mu,\bm{r},\bm{s},I)=D(\mu,\mu,\bm{r},\bm{s},I)=1.$
Proof.
By definition we have $M(\mu,\mu,\bm{r},\bm{s},I)=1$. The other equation $D(\mu,\mu,\bm{r},\bm{s},I)=1$ can easily be proved by observing that the matrix in the definition $D(\mu,\mu,\bm{r},\bm{s},I)=\det(h_{\mu,\mu}^{\bm{r},\bm{s}}(I;i,j))$ is upper unitriangular. ∎
The following two lemmas give the same recursions for $M(\lambda,\mu,\bm{r},\bm{s},I)$ and $D(\lambda,\mu,\bm{r},\bm{s},I)$.
Lemma 4.17.
Let $\lambda\in\operatorname{DPar}_{n},\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. Suppose that $11\leq t\leq n-$ 1 is an integer such that $\mu_{t}\geq\lambda_{t+1}$. Let $\phi_{\bm{\beta}}$ be the map defined by shifting $\beta_{i}$ to $\beta_{i+1}$ for $i\geq 1$. Then for $F\in\{D,M\}$, we have
$F(\lambda,\mu,\bm{r},\bm{s},I)=F(\lambda^{(1)},\mu^{(1)},\bm{r}^{(1)},\bm{s}^{(1)},I^{(1)})\phi_{\bm{\beta}}^{t}(F(\lambda^{(2)},\mu^{(2)},\bm{r}^{(2)},\bm{s}^{(2)},I^{(2)})),$
where $\gamma^{(1)}=(\gamma_{1},\dots,\gamma_{t})$ and $\gamma^{(2)}=(\gamma_{t+1},\dots,\gamma_{n})$ for each $\gamma\in\{\lambda,\mu,\bm{r},\bm{s}\}$, and $I^{(1)}=I\cap\{1,2,\dots,t\}$ and $I^{(2)}=I\cap\{t+1,t+2,\dots,n\}$.
Proof.
If $F=M$, the statement follow immediately from the definition of $M(\lambda,\mu,\bm{r},\bm{s},I)$. For the case $F=D$, since
$D(\lambda^{(1)},\mu^{(1)},\bm{r}^{(1)},\bm{s}^{(1)},I^{(1)})$ $=\det\left(h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)\right)_{i,j=1}^{t},$
$\phi_{\bm{\beta}}^{t}(D(\lambda^{(2)},\mu^{(2)},\bm{r}^{(2)},\bm{s}^{(2)},I^{(2)}))$ $=\det\left(h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)\right)_{i,j=t+1}^{n},$
it suffices to show that $h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)=0$ for all $t+1\leq i\leq n$ and $1\leq j\leq t$. Suppose $1\leq j\leq t$. If $i=t+1$, then
$\lambda_{i}-i-\mu_{j}+j\leq\lambda_{t+1}-(t+1)-\mu_{t}+t\leq-1,$
and therefore $h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)=0$. If $t+2\leq i\leq n$, since $\lambda_{b}\leq\lambda_{a}+1$ for any $1\leq a<b\leq n$, we have
$\lambda_{i}-i-\mu_{j}+j\leq(\lambda_{t+1}+1)-(t+2)-\mu_{t}+t\leq-1.$
Thus $h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)=0$ and we obtain the desired identity. ∎
Lemma 4.18.
Let $\lambda\in\operatorname{DPar}_{n}$ be a dented partition with the minimal cell $(k,\lambda_{k})$ and let $\mu\in\operatorname{Par}_{n}$ be a partition with $\mu\subseteq\lambda\setminus\{(k,\lambda_{k})\}$. Let $\bm{r},\bm{s}\in\operatorname{RPar}_{n}$ with $\bm{r}\leq\bm{s}$. Suppose that $I=\{1,2,\dots,p\}$ or $I=\{1,2,\dots,p\}\setminus\{k\}$ for some integer $k\leq p\leq\ell(\lambda)$ satisfying
$\lambda_{1}+1=\dots=\lambda_{k-1}+1=\lambda_{k}=\lambda_{k+1}=\dots=\lambda_{p}.$
If $\lambda$ is not a partition and $s_{k-1}=s_{k}$, then for $F\in\{D,M\}$,
$F(\lambda,\mu,\bm{r},\bm{s},I)=(\beta_{k-1}-\chi(k\in I)\alpha_{\lambda_{k}})F(\lambda-\epsilon_{k},\mu,\bm{r},\bm{s},I\cup\{k\}).$
If $\lambda$ is a partition or $s_{k-1}<s_{k}$, then for $F\in\{D,M\}$,
$F(\lambda,\mu,\bm{r},\bm{s},I)=(x_{s_{k}}-\chi(k\in I)\alpha_{\lambda_{k}})F(\lambda-\epsilon_{k},\mu,\bm{r},\bm{s},I\cup\{k\})+F(\lambda,\mu,\bm{r},\bm{s}-\epsilon_{k},I\setminus\{k\}).$
Proof.
If $F=M$, then the two statements follow immediately from the definition of $M(\lambda,\mu,\bm{r},\bm{s},I)$. For the case $F=D$, let
$H=\left(h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)\right)_{i,j=1}^{n}.$
For the first statement suppose that $\lambda$ is not a partition and $s_{k-1}=s_{k}$. We claim that for all $1\leq j\leq n$,
(4.7) $h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)=h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k-1,j)+(\beta_{k-1}-\chi(k\in I)\alpha_{\lambda_{k}})h_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j).$
Note that by subtracting the $(k-1)$st row from the $k$th row in $H$, the claim implies the first statement. Thus it suffices to prove (4.7).
For brevity, let
(4.8) $m=\lambda_{k}-\mu_{j}+j-k,\qquad Z=X_{[r_{j},s_{k}]}-A_{\lambda_{k}-1}+A_{\mu_{j}}+B_{k-1}-B_{j-1}.$
Then
$h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)$ $=\chi(r_{j}\leq s_{k})h_{\lambda_{k}-\mu_{j}+j-k}[X_{[r_{j},s_{k}]}-A_{\lambda_{k}-1+\chi(k\in I)}+A_{\mu_{j}}+B_{k-1}-B_{j-1}]$
(4.9) $=\chi(r_{j}\leq s_{k})h_{m}[Z-\chi(k\in I)\alpha_{\lambda_{k}}],$
where the second equality follows from $A_{\lambda_{k}-1+\chi(k\in I)}=A_{\lambda_{k}-1}+\chi(k\in I)\alpha_{\lambda_{k}}$. By Lemma 2.2,
$h_{m}[Z-\chi(k\in I)\alpha_{\lambda_{k}}]$ $=h_{m}[Z]-\chi(k\in I)\alpha_{\lambda_{k}}h_{m-1}[Z]$
(4.10) $=h_{m}[Z-\beta_{k-1}]+\beta_{k-1}h_{m-1}[Z]-\chi(k\in I)\alpha_{\lambda_{k}}h_{m-1}[Z].$
On the other hand, since $\lambda_{k-1}=\lambda_{k}-1$ and $k-1\in I$,
$h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k-1,j)$ $=\chi(r_{j}\leq s_{k-1})h_{\lambda_{k-1}-\mu_{j}+j-(k-1)}[X_{[r_{j},s_{k-1}]}-A_{\lambda_{k-1}}+A_{\mu_{j}}+B_{k-2}-B_{j-1}]$
$=\chi(r_{j}\leq s_{k})h_{\lambda_{k}-\mu_{j}+j-k}[X_{[r_{j},s_{k}]}-A_{\lambda_{k}-1}+A_{\mu_{j}}+B_{k-2}-B_{j-1}]$
(4.11) $=\chi(r_{j}\leq s_{k})h_{m}[Z-\beta_{k-1}].$
Since $h_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j)=\chi(r_{j}\leq s_{k})h_{m-1}[Z]$, (4.9), (4.10), and (4.11) give the claim (4.7). This shows the first statement.
For the second statement suppose that $\lambda$ is a partition or $s_{k-1}<s_{k}$. We claim that for all $1\leq j\leq n$,
(4.12) $h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)=h_{\lambda,\mu}^{\bm{r},\bm{s}-\epsilon_{k}}(I\setminus\{k\};k,j)+(x_{s_{k}}-\chi(k\in I)\alpha_{\lambda_{k}})h_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j).$
Note that by the linearity of the determinant in the $k$th row of $H$, the claim implies the second statement. Thus it suffices to prove (4.12).
We again use the notation $m$ and $Z$ in (4.8). By the same computation as in (4.10) with $\beta_{k-1}$ replaced by $x_{s_{k}}$, we obtain
$h_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)$ $=\chi(r_{j}\leq s_{k})h_{m}[Z-x_{s_{k}}]+\chi(r_{j}\leq s_{k})(x_{s_{k}}-\chi(k\in I)\alpha_{\lambda_{k}})h_{m-1}[Z]$
$=\chi(r_{j}\leq s_{k})h_{m}[Z-x_{s_{k}}]+(x_{s_{k}}-\chi(k\in I)\alpha_{\lambda_{k}})h_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j).$
Since $h_{\lambda,\mu}^{\bm{r},\bm{s}-\epsilon_{k}}(I\setminus\{k\};k,j)=\chi(r_{j}\leq s_{k}-1)h_{m}[Z-x_{s_{k}}]$, by the above equation, in order to prove (4.12) it suffices show that
$\chi(r_{j}\leq s_{k})h_{m}[Z-x_{s_{k}}]=\chi(r_{j}\leq s_{k}-1)h_{m}[Z-x_{s_{k}}].$
If $r_{j}\neq s_{k}$, then $\chi(r_{j}\leq s_{k})=\chi(r_{j}\leq s_{k}-1)$ and we are done. Thus we may assume $r_{j}=s_{k}$. In this case we must show that $h_{m}[Z-x_{s_{k}}]=0$, i.e.,
(4.13) $h_{\lambda_{k}-\mu_{j}+j-k}[-A_{\lambda_{k}-1}+A_{\mu_{j}}+B_{k-1}-B_{j-1}]=0.$
To prove (4.13) we consider the two cases $j<k$ and $j\geq k$
Suppose $j<k$. In this case we have $k\geq 2$. Since $r_{k-1}\leq s_{k-1}<s_{k}=r_{j}$ and $\bm{r}\in\operatorname{RPar}_{n}$, we have $k-1<j$, which is a contradiction to $j<k$.
We now suppose $j\geq k$. If $k=1$, then by assumption we have $\lambda_{k}-1\geq\mu_{k}\geq\mu_{j}$. If $k\geq 2$, since $\lambda$ is a dented partition and $\mu$ is a partition contained in $\lambda$, we have $\lambda_{k}-1=\lambda_{k-1}\geq\mu_{k-1}\geq\mu_{j}$. Therefore we always have $\lambda_{k}-1\geq\mu_{j}$ and the left side of (4.13) is equal to
$(-1)^{\lambda_{k}-\mu_{j}+j-k}e_{\lambda_{k}-\mu_{j}+j-k}[A_{\lambda_{k}-1}-A_{\mu_{j}}-B_{k-1}+B_{j-1}]\\ =(-1)^{\lambda_{k}-\mu_{j}+j-k}e_{\lambda_{k}-\mu_{j}+j-k}(\alpha_{\mu_{j}+1},\dots,\alpha_{\lambda_{k}-1},\beta_{k},\dots,\beta_{j-1}),$
which is equal to 00 because the number $\lambda_{k}-1-\mu_{j}+(j-1)-(k-1)$ of variables is less than $\lambda_{k}-\mu_{j}+j-k$. ∎
Now we are ready to prove Proposition 4.14.
Proof of Proposition 4.14.
We proceed by induction on $(M,N)$, where $M:=\sum_{i=1}^{n}(s_{i}-r_{i})$ and $N:=|{\lambda/\mu}|$. There are two base cases: the case $M<0$ and the case $N=0$, which can be easily obtained from Lemmas 4.15 and 4.16.
For the inductive step, suppose that $M\geq 0,N\geq 1$ and the result holds for all pairs $(M^{\prime},N^{\prime})$ with $M^{\prime}<M$ or $N^{\prime}<N$. Now consider a tuple $(\lambda,\mu,\bm{r},\bm{s},I)$ satisfying the conditions in Proposition 4.14 such that $\sum_{i=1}^{n}(s_{i}-r_{i})=M$ and $|{\lambda/\mu}|=N$. If $\bm{r}\not\leq\bm{s}$, then the result holds by Lemma 4.15. Therefore we may assume $\bm{r}\leq\bm{s}$. We consider the following two cases.
Case 1: Suppose that $\mu_{t}<\lambda_{t+1}$ for all $11\leq t\leq n-$ 1. By the assumptions on $\bm{r}$ and $\bm{s}$, in this case we have $\bm{r},\bm{s}\in\operatorname{RPar}_{n}$. Let $(k,\lambda_{k})$ be the minimal cell of $\lambda$. Since $\mu_{k}<\lambda_{k+1}\leq\lambda_{k}$, we have $(k,\lambda_{k})\in{\lambda/\mu}$. Hence by Lemma 4.18 we can decrease either $M=\sum(s_{i}-r_{i})$ or $N=|{\lambda/\mu}|$. Thus this case is obtained by the induction hypothesis.
Case 2: Suppose that an integer $11\leq t\leq n-$ 1 satisfies $\mu_{t}\geq\lambda_{t+1}$. By Lemma 4.17, it suffices to show that
$M(\lambda^{(1)},\mu^{(1)},\bm{r}^{(1)},\bm{s}^{(1)},I^{(1)})$ $=D(\lambda^{(1)},\mu^{(1)},\bm{r}^{(1)},\bm{s}^{(1)},I^{(1)})\mbox{ and }$
$M(\lambda^{(2)},\mu^{(2)},\bm{r}^{(2)},\bm{s}^{(2)},I^{(2)})$ $=D(\lambda^{(2)},\mu^{(2)},\bm{r}^{(2)},\bm{s}^{(2)},I^{(2)}),$
where we follow the notations in Lemma 4.17. In this way we can reduce the size of $n$ without increasing $M$ or $N$. Observe that if $n=1$ we must have Case 1 above. Therefore by repeating this process we can divide the problem of proving the result for the tuple $(\lambda,\mu,\bm{r},\bm{s},I)$ into those for some tuples $(\lambda^{*},\mu^{*},\bm{r}^{*},\bm{s}^{*},I^{*})$ with $M^{*}=\sum(s^{*}_{i}-r^{*}_{i})\leq M$ and $N^{*}=|\lambda^{*}/\mu^{*}|\leq N$ that are all in Case 1. Then this case is obtained by the induction hypothesis and by Case 1. ∎
4.2. Proof of (4.3): column-flagged dual Grothendieck polynomials
In this subsection we prove the second assertion (4.3) of Theorem 4.5. Our strategy is similar to that in the previous subsection. We first show that (4.3) is equivalent to
(4.14) $g_{\lambda^{\prime}/\mu^{\prime}}^{\operatorname{col}(\bm{r},\bm{s})}(\bm{x};\bm{\alpha},\bm{\beta})\\ =\det\left(\chi(r_{j}\leq s_{i})e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1}-B_{\mu_{j}}]\right)_{i,j=1}^{n}.$
We then find a generalization of this formula and prove it using recursions.
The following lemma shows that (4.3) and (4.14) are equivalent.
Lemma 4.19.
Let $\lambda,\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$ and $\bm{r}\leq\bm{s}$ such that $r_{i}\leq r_{i+1}$ and $s_{i}\leq s_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$. Then
$\det\left(\chi(r_{j}\leq s_{i})e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1}-B_{\mu_{j}}]\right)_{i,j=1}^{n}\\ =\det\left(e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1}-B_{\mu_{j}}]\right)_{i,j=1}^{n}.$
Proof.
This can be proved by the same argument in the proof of Lemma 4.10. The only difference is the following analogue of (4.4): if $i<j$ and $\mu_{j}<\lambda_{i}$, then we have
$e_{\lambda_{i}-\mu_{j}-i+j}[-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1}-B_{\mu_{j}}]=e_{\lambda_{i}-\mu_{j}-i+j}(\alpha_{i},\dots,\alpha_{j-1},\beta_{\mu_{j}+1},\dots,\beta_{\lambda_{i}-1})=0.$
Then the argument in the proof of Lemma 4.10 with this fact completes the proof. ∎
Our next step is to find a generalization of (4.14). We first reformulate (4.14). To do this we introduce two different versions of marked RPPs.
Definition 4.20.
A right-marked RPP of shape ${\lambda/\mu}$ is an RPP $T$ of shape ${\lambda/\mu}$ in which every entry $T(i,j)$ with $T(i,j)=T(i,j-1)$ may be marked. We denote by ${\operatorname{RMRPP}}({\lambda/\mu})$ the set of right-marked RPPs of shape ${\lambda/\mu}$. For $T\in{\operatorname{RMRPP}}({\lambda/\mu})$, define
$\operatorname{wt}_{R}(T)=\prod_{(i,j)\in{\lambda/\mu}}\operatorname{wt}_{R}(T(i,j)),$
where
$\operatorname{wt}_{R}(T(i,j))=\begin{cases}-\alpha_{j-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{i}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i+1,j)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i+1,j)$.}\end{cases}$
See Figure 6.
[…]
Figure 6. An example of $T\in{\operatorname{RMRPP}}({\lambda/\mu})$, where $\lambda=(6,5,3,3)$ and $\mu=(2,1,1)$. We have the weight $\operatorname{wt}(T)=x_{1}^{2}x_{2}x_{3}^{2}x_{4}x_{5}(-\alpha_{1})(-\alpha_{2})^{2}(-\alpha_{5})\beta_{1}\beta_{2}$.
Definition 4.21.
A bottom-marked RPP of shape ${\lambda/\mu}$ is an RPP $T$ of shape ${\lambda/\mu}$ in which every entry $T(i,j)$ with $T(i,j)=T(i-1,j)$ may be marked. We denote by ${\operatorname{BMRPP}}({\lambda/\mu})$ the set of bottom-marked RPPs of shape ${\lambda/\mu}$. For $T\in{\operatorname{BMRPP}}({\lambda/\mu})$, define
$\operatorname{wt}_{B}(T)=\prod_{(i,j)\in{\lambda/\mu}}\operatorname{wt}_{B}(T(i,j)),$
where
$\operatorname{wt}_{B}(T(i,j))=\begin{cases}-\alpha_{i-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{j}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i,j+1)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i,j+1)$.}\end{cases}$
One can easily see that bottom-marked RPPs are just the transposes of right-marked RPPs, and their weights coincide, that is, for $T\in{\operatorname{RMRPP}}(\lambda/\mu)$, $\operatorname{wt}_{R}(T)=\operatorname{wt}_{B}(T^{\prime})$.
For an RPP $T_{0}$ of shape ${\lambda/\mu}$, let ${\operatorname{MRPP}}(T_{0})$ (resp. ${\operatorname{RMRPP}}(T_{0})$) be the set of left-marked (resp. right-marked) RPPs $T$ whose underlying RPP is $T_{0}$. To avoid confusion, in this section we also write $\operatorname{wt}_{L}$ for the weight function on left-marked RPPs given in Definition 4.1.
Proposition 4.22.
For an RPP $T_{0}$ of shape ${\lambda/\mu}$,
$\sum_{T\in{\operatorname{MRPP}}(T_{0})}\operatorname{wt}_{L}(T)=\sum_{T\in{\operatorname{RMRPP}}(T_{0})}\operatorname{wt}_{R}(T).$
Proof.
For $k\geq 1$, let $T_{0}[k]$ be the RPP obtained from $T_{0}$ by taking the cells filled with $k$. Then, by definition,
$\sum_{T\in{\operatorname{MRPP}}(T_{0})}\operatorname{wt}_{L}(T)=\prod_{k\geq 1}\sum_{T\in{\operatorname{MRPP}}(T_{0}[k])}\operatorname{wt}_{L}(T)$
and
$\sum_{T\in{\operatorname{RMRPP}}(T_{0})}\operatorname{wt}_{R}(T)=\prod_{k\geq 1}\sum_{T\in{\operatorname{RMRPP}}(T_{0}[k])}\operatorname{wt}_{R}(T).$
Therefore, without loss of generality we may assume that every entry in $T_{0}$ is $k$ for some fixed $k\geq 1$. Then, for a left-marked RPP $T\in{\operatorname{MRPP}}(T_{0})$, we have
(4.15) $\operatorname{wt}_{L}(T)=x_{k}^{m_{L}}\prod_{j\geq 1}(-\alpha_{j})^{a_{L,j}}\prod_{i\geq 1}\beta_{i}^{b_{L,i}},$
where $m_{L}$ is the number of cells $u$ with $k$ in $T$ such that the cell above $u$ is not in ${\lambda/\mu}$, $a_{L,j}$ is the number of cells with $k^{*}$ in the $j$th column of $T$, and $b_{L,i}$ is the number of cells $u$ with $k$ in the $(i+1)$st row of $T$ such that the cell above $u$ is in ${\lambda/\mu}$. Similarly, for a right-marked RPP $T\in{\operatorname{RMRPP}}(T_{0})$,
(4.16) $\operatorname{wt}_{R}(T)=x_{k}^{m_{R}}\prod_{j\geq 1}(-\alpha_{j})^{a_{R,j}}\prod_{i\geq 1}\beta_{i}^{b_{R,i}},$
where $m_{R}$ is the number of cells $u$ with $k$ in $T$ such that the cell below $u$ is not in ${\lambda/\mu}$, $a_{R,j}$ is the number of cells with $k^{*}$ in the $(j+1)$st column of $T$, and $b_{R,i}$ is the number of cells $u$ with $k$ in the $i$th row of $T$ such that the cell below $u$ is in ${\lambda/\mu}$.
For a left-marked RPP $T\in{\operatorname{MRPP}}(T_{0})$, let $\varphi(T)$ be the right-marked RPP in ${\operatorname{RMRPP}}(T_{0})$ such that, assuming $(i,j),(i,j-1)\in{\lambda/\mu}$, the cell $(i,j)$ is marked in $\varphi(T)$ if and only if the cell $(i+1,j-1)$ is marked in $T$ if $(i+1,j-1)\in\lambda$, and the cell $(\mu^{\prime}_{j-1}+1,j-1)$ is marked in $T$ otherwise. In other words, the map $\varphi$ first moves each marking to its upper cell with the identification $(\lambda^{\prime}_{j},j)=(\mu^{\prime}_{j},j)$ of the last cells in each column of $\lambda$ and $\mu$, and then moves to the right cell. Then using (4.15) and (4.16), one can directly check that $\operatorname{wt}_{L}(T)=\operatorname{wt}_{R}(\varphi(T))$. Clearly, the map $\varphi$ is bijective, and the proof follows. See Figure 7. ∎
[…]
Figure 7. An example of the map $\varphi$ sending $T$ to $\varphi(T)$ for $\lambda=(5,5,4,4,4,3),\mu=(3,2,2,1,1)$, and $T$ with all cells filled with $k$ or $k^{*}$. The arrows on the left diagram show how the marked integers and unmarked integers move from column 3 to column 4.
For $\bm{r},\bm{s}\in\mathbb{P}^{n}$, we denote by ${\operatorname{RMRPP}}^{\operatorname{col}(\bm{r},\bm{s})}(\lambda^{\prime}/\mu^{\prime})$ the set of right-marked RPPs $T$ of shape $\lambda^{\prime}/\mu^{\prime}$ such that for each $(i,j)\in\lambda^{\prime}/\mu^{\prime}$, $r_{j}\leq T(i,j)\leq s_{j}$. According to Proposition 4.22, we can restate (4.14) as
(4.17) $\sum_{T\in{\operatorname{RMRPP}}^{\operatorname{col}(\bm{r},\bm{s})}(\lambda^{\prime}/\mu^{\prime})}\operatorname{wt}_{R}(T)\\ =\det\left(\chi(r_{j}\leq s_{i})e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1}-B_{\mu_{j}}]\right)_{i,j=1}^{n}.$
To apply the inductive arguments in the previous section, we consider the transposes of right-marked RPPs so that the column bound conditions $r_{j}\leq T(i,j)\leq s_{j}$ are changed into the row bound conditions $r_{i}\leq T(i,j)\leq s_{i}$. Since the transpose of a right-marked RPP is a bottom-marked RPP, we define ${\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}(\lambda/\mu)$ similarly to ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}(\lambda/\mu)$, and then we can restate (4.17) as follows:
(4.18) $\sum_{T\in{\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}(\lambda/\mu)}\operatorname{wt}_{B}(T)\\ =\det\left(\chi(r_{j}\leq s_{i})e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}]}-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1}-B_{\mu_{j}}]\right)_{i,j=1}^{n}.$
We also introduce a notion similar to ${\operatorname{MRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}(\lambda/\mu)$.
Definition 4.23.
Let $\lambda$ be a dented partition and $\mu$ a partition with $\mu\subseteq\lambda$. Let $\bm{r},\bm{s}\in\mathbb{P}^{n}$ and $I\subseteq[n]$. If $\bm{r}\not\leq\bm{s}$, we define ${\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}(\lambda/\mu)=\emptyset$. If $\bm{r}\leq\bm{s}$, we define ${\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$ to be the set of RPPs $T$ of shape ${\lambda/\mu}$ in which every entry $T(i,j)$ with $T(i,j)=T(i-1,j)$ may be marked. For each $i\in[n]$, we define
$T(i,\lambda_{i}+1)=\begin{cases}\infty&\mbox{if $i\not\in I$},\\ s_{i}&\mbox{if $i\in I$.}\end{cases}$
For $T\in{\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})$, we define
$\operatorname{wt}_{B}(T)=\prod_{(i,j)\in{\lambda/\mu}}\operatorname{wt}_{B}(T(i,j)),$
where
$\operatorname{wt}_{B}(T(i,j))=\begin{cases}-\alpha_{i-1}&\mbox{if$T(i,j)$ is marked,}\\ \beta_{j}&\mbox{if $T(i,j)$ is not marked and $T(i,j)=T(i,j+1)$,}\\ x_{T(i,j)}&\mbox{if $T(i,j)$ is not marked and $T(i,j)\neq T(i,j+1)$.}\end{cases}$
Finally we state a generalization of (4.18), which has been shown to be equivalent to the Jacobi–Trudi-like identity in (4.3).
Proposition 4.24.
Let $\lambda\in\operatorname{DPar}_{n},\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$ such that $r_{i}\leq r_{i+1}$ and $s_{i}\leq s_{i+1}$ whenever $\mu_{i}<\lambda_{i+1}$. Suppose that $(k,\lambda_{k})$ is the minimal cell of $\lambda$ and $I=\{1,2,\dots,p\}$ or $I=\{1,2,\dots,p\}\setminus\{k\}$ for some integer $k\leq p\leq\ell(\lambda)$ satisfying
$\lambda_{1}+1=\dots=\lambda_{k-1}+1=\lambda_{k}=\lambda_{k+1}=\dots=\lambda_{p}.$
Then
(4.19) $\sum_{T\in{\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})}\operatorname{wt}_{B}(T)=\det\left(e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)\right)_{i,j=1}^{n},$
where
$e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)=\chi(r_{j}\leq s_{i})e_{\lambda_{i}-\mu_{j}-i+j}[X_{[r_{j},s_{i}-\chi(i\in I)]}-A_{i-1}+A_{j-1}+B_{\lambda_{i}-1+\chi(i\in I)}-B_{\mu_{j}}].$
Observe that (4.18) is the special case of Proposition 4.24 when $\lambda$ is a partition and $I=\emptyset$. Similar to the previous subsection we prove Proposition 4.24 by showing that both sides of the equation have the same recursions and initial conditions. Let
$M^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)$ $=\sum_{T\in{\operatorname{BMRPP}}^{\operatorname{row}(\bm{r},\bm{s})}_{I}({\lambda/\mu})}\operatorname{wt}_{B}(T),$
$D^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)$ $=\det\left(e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;i,j)\right)_{i,j=1}^{n}.$
The next two lemmas give the same initial conditions for $M^{\prime}$ and $D^{\prime}$. These lemmas can be proved by the same arguments in the proofs of Lemmas 4.15 and 4.16, respectively.
Lemma 4.25.
Let $\lambda\in\operatorname{DPar}_{n},\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\operatorname{RPar}_{n}$ with $\mu\subseteq\lambda$. If $\bm{r}\not\leq\bm{s}$, then
$M^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)=D^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)=0.$
Lemma 4.26.
Let $\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$. If $\bm{r}\leq\bm{s}$, then
$M^{\prime}(\mu,\mu,\bm{r},\bm{s},I)=D^{\prime}(\mu,\mu,\bm{r},\bm{s},I)=1.$
The following two lemmas give the same recursions for $M^{\prime}$ and $D^{\prime}$.
Lemma 4.27.
Let $\lambda\in\operatorname{DPar}_{n},\mu\in\operatorname{Par}_{n},\bm{r},\bm{s}\in\mathbb{P}^{n}$ with $\mu\subseteq\lambda$. Suppose that $11\leq t\leq n-$ 1 is an integer such that $\mu_{t}\geq\lambda_{t+1}$. Let $\phi_{\bm{\alpha}}$ be the map defined by shifting $\alpha_{i}$ to $\alpha_{i+1}$ for $i\geq 1$, then for $F\in\{M,D\}$, we have
$F^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)=F^{\prime}(\lambda^{(1)},\mu^{(1)},\bm{r}^{(1)},\bm{s}^{(1)},I^{(1)})\phi_{\bm{\alpha}}^{t}(F^{\prime}(\lambda^{(2)},\mu^{(2)},\bm{r}^{(2)},\bm{s}^{(2)},I^{(2)})),$
where $\gamma^{(1)}=(\gamma_{1},\dots,\gamma_{t})$ and $\gamma^{(2)}=(\gamma_{t+1},\dots,\gamma_{n})$ for each $\gamma\in\{\lambda,\mu,\bm{r},\bm{s}\}$, and $I^{(1)}=I\cap\{1,2,\dots,t\}$ and $I^{(2)}=I\cap\{t+1,t+2,\dots,n\}$.
Proof.
The proof follows from the same argument of the proof of Lemma 4.17. ∎
Lemma 4.28.
Let $\lambda\in\operatorname{DPar}_{n}$ be a dented partition with the minimal cell $(k,\lambda_{k})$ and let $\mu\in\operatorname{Par}_{n}$ be a partition with $\mu\subseteq\lambda\setminus\{(k,\lambda_{k})\}$. Let $\bm{r},\bm{s}\in\operatorname{RPar}_{n}$ with $\bm{r}\leq\bm{s}$. Suppose that $I=\{1,2,\dots,p\}$ or $I=\{1,2,\dots,p\}\setminus\{k\}$ for some integer $k\leq p\leq\ell(\lambda)$ satisfying
$\lambda_{1}+1=\dots=\lambda_{k-1}+1=\lambda_{k}=\lambda_{k+1}=\dots=\lambda_{p}.$
If $\lambda$ is not a partition and $s_{k-1}=s_{k}$, then for $F\in\{M,D\}$,
$F^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)=(x_{s_{k}}-\alpha_{k-1}+\chi(k\in I)\left(\beta_{\lambda_{k}}-x_{s_{k}}\right))F^{\prime}(\lambda-\epsilon_{k},\mu,\bm{r},\bm{s},I\cup\{k\}).$
If $\lambda$ is a partition or $s_{k-1}<s_{k}$, then for $F\in\{M,D\}$,
$F^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)=(x_{s_{k}}+\chi(k\in I)\left(\beta_{\lambda_{k}}-x_{s_{k}}\right))F^{\prime}(\lambda-\epsilon_{k},\mu,\bm{r},\bm{s},I\cup\{k\})\\ +F^{\prime}(\lambda,\mu,\bm{r},\bm{s}-\epsilon_{k},I\setminus\{k\}).$
Proof.
The case $F=M$ is immediate from the definition of $M^{\prime}(\lambda,\mu,\bm{r},\bm{s},I)$, so we only consider the case $F=D$.
For the first assertion, suppose that $\lambda$ is not a partition and $s_{k-1}=s_{k}$. Following the argument in the proof of Lemma 4.18, we only need to prove that for $1\leq j\leq n$,
(4.20) $e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)=e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k-1,j)+(x_{s_{k}}-\alpha_{k-1}+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}}))e_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j),$
which is an analogue of (4.7). Let
$m=\lambda_{k}-\mu_{j}+j-k\quad\mbox{and}\quad Z=X_{[r_{j},s_{k}]}-A_{k-1}+A_{j-1}+B_{\lambda_{k}-1}-B_{\mu_{j}},$
then, since $s_{k-1}=s_{k}$ and $\lambda_{k-1}=\lambda_{k}-1$, we can rewrite each term in (4.20) as
$e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)$ $=\chi(r_{j}\leq s_{k})e_{m}[Z+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}})],$
$e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k-1,j)$ $=\chi(r_{j}\leq s_{k})e_{m}[Z-x_{s_{k}}+\alpha_{k-1}],$
$e_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j)$ $=\chi(r_{j}\leq s_{k})e_{m-1}[Z-x_{s_{k}}].$
By iterating Lemma 2.3,
$e_{m}[Z+\beta_{\lambda_{k}}-x_{s_{k}}]=e_{m}[Z-x_{s_{k}}]+\beta_{\lambda_{k}}e_{m-1}[Z-x_{s_{k}}]=e_{m}[Z]+(\beta_{\lambda_{k}}-x_{s_{k}})e_{m-1}[Z-x_{s_{k}}],$
and thus we can write
(4.21) $e_{m}[Z+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}})]=e_{m}[Z]+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}})e_{m-1}[Z-x_{s_{k}}].$
By iterating Lemma 2.3, we similarly obtain
$e_{m}[Z]=e_{m}[Z-x_{s_{k}}+\alpha_{k-1}]+(x_{s_{k}}-\alpha_{k-1})e_{m-1}[Z-x_{s_{k}}].$
By (4.21) and the above equation, $e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)$ is equal to
$\chi(r_{j}\leq s_{k})\left(e_{m}[Z]+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}})e_{m-1}[Z-x_{s_{k}}]\right)$
$=\chi(r_{j}\leq s_{k})\left(e_{m}[Z-x_{s_{k}}+\alpha_{k-1}]+(x_{s_{k}}-\alpha_{k-1}+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}}))e_{m-1}[Z-x_{s_{k}}]\right)$
$=e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k-1,j)+(x_{s_{k}}-\alpha_{k-1}+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}}))e_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j),$
which shows the first assertion.
For the second assertion, suppose that $\lambda$ is a partition or $s_{k-1}<s_{k}$. Similar to the first case, it suffices to show that for $1\leq j\leq n$,
(4.22) $e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)=e_{\lambda,\mu}^{\bm{r},\bm{s}-\epsilon_{k}}(I\setminus\{k\};k,j)+(x_{s_{k}}+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}}))e_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j).$
Using the terms $Z$ and $m$ defined above, we can write
$e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)$ $=\chi(r_{j}\leq s_{k})e_{m}[Z+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}})],$
$e_{\lambda,\mu}^{\bm{r},\bm{s}-\epsilon_{k}}(I\setminus\{k\};k,j)$ $=\chi(r_{j}\leq s_{k}-1)e_{m}[Z-x_{s_{k}}],$
$e_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j)$ $=\chi(r_{j}\leq s_{k})e_{m-1}[Z-x_{s_{k}}].$
Then by (4.21) and Lemma 2.3, we obtain
$e_{\lambda,\mu}^{\bm{r},\bm{s}}(I;k,j)$ $=\chi(r_{j}\leq s_{k})\left(e_{m}[Z]+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}})e_{m}[Z-x_{s_{k}}]\right)$
$=\chi(r_{j}\leq s_{k})\left(e_{m}[Z-x_{s_{k}}]+(x_{s_{k}}+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}}))e_{m-1}[Z-x_{s_{k}}]\right)$
$=\chi(r_{j}\leq s_{k})e_{m}[Z-x_{s_{k}}]+(x_{s_{k}}+\chi(k\in I)(\beta_{\lambda_{k}}-x_{s_{k}}))e_{\lambda-\epsilon_{k},\mu}^{\bm{r},\bm{s}}(I\cup\{k\};k,j).$
Hence, to prove (4.22) it is enough to show
$\chi(r_{j}\leq s_{k}-1)e_{m}[Z-x_{s_{k}}]=\chi(r_{j}\leq s_{k})e_{m}[Z-x_{s_{k}}].$
Obviously, if $r_{j}\neq s_{k}$, then the both sides are the same. If $r_{j}=s_{k}$, the above equation is equivalent to
$e_{\lambda_{k}-\mu_{j}+j-k}[-A_{k-1}+A_{j-1}+B_{\lambda_{k}-1}-B_{\mu_{j}}]=0.$
This can be prove similarly as (4.13), and thus we obtain (4.22). ∎
Finally we are ready to prove Proposition 4.24. This can be done by the same arguments in the proof of Proposition 4.14 with Lemmas 4.25, 4.26, 4.28, and 4.27 in place of Lemmas 4.15, 4.16, 4.18 and 4.17. We omit the details.
Acknowledgments
The authors are grateful to Travis Scrimshaw for many helpful comments and fruitful discussions. They also thank Darij Grinberg for useful comments.